### v2

# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

Abstract
Understanding how user preference evolves over time is a fundamental challenge central to modern digital ecosystems, for which Large Language Models (LLMs) are an increasingly prominent and popular approach due to their ability to comprehend the rich semantic context within behavioral data. A common practice is to use LLMs to predict a user’s next action by directly generating a ranked list of preferred items. Although effective for short-term prediction, the end-to-end generation paradigm inherently limits personalization. Its opaque decision-making process obscures holistic user profiling and exacerbates popularity bias. To address these limitations, we propose Preference Evolution Tracking (PET), a framework that reframes the task as inferring a dynamic probability distribution over a stable and interpretable lattice of preference clusters. By applying logit-probing and generative classification techniques, PET infers a user’s preference as a probability distribution, enabling transparent preference learning. On public benchmarks (Yelp, MovieLens), PET improves ranking quality by up to $40\%$ in NDCG over direct generation baselines. On a large-scale, real-world dataset from a short-video platform, it excels at ranking long-tail contents, significantly outperforming a SOTA production model by $7$ times in the NDCG score. Ultimately, PET transforms the user profile model from direct preference list generation to a transparent distributional preference mapping, paving the way for more explainable, fair, and diverse personalization systems.
1 Introduction
Understanding user interests in a transparent and interpretable manner is a central challenge in building trusted and adaptive digital ecosystems. The advent of Large Language Models (LLMs) presents a transformative opportunity in this domain. With their profound ability to comprehend context and semantics within unstructured user behavioral histories – from product reviews to content consumption sequences – LLMs promise to move beyond simple pattern matching to a deeper, more holistic understanding of user intent. The ultimate goal is not merely to predict the next action of users, but to create transparent and dynamic models of user interest that can be trusted, analyzed, and fairly acted upon.
The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked outputs (wang2025user; wang2024llms; ngo2024recgpt; yu2024icpl; deng2025onerec), a method proven effective for short-horizon, top-$k$ recommendation tasks. While effective for short-horizon prediction and identifying popular preferences, this direct-to-ranking paradigm leaves important gaps. First, it lacks interpretability: although LLMs encode rich internal states, these models do not expose an explicit or structured representation of the user’s overall preferences – making them difficult to audit or reuse in downstream tasks. Second, the focus on a small $k$, combined with popularity bias, often leads to the over-representation of mainstream content – neglecting the long-tail interests that define a user’s niche tastes and are critical for achieving holistic, unbiased personalization.
These limitations have significant consequences for algorithmic decision-making in user-oriented applications. The lack of an explicit user model undermines system transparency, making it difficult to govern and audit for fairness and bias. Additionally, the over-reliance on popular items not only distorts model predictions but also reinforces popularity during training – suppressing the model’s ability to learn users’ nuanced, long-tail preferences. This feedback loop amplifies the Matthew Effect (i.e., the rich get richer), reinforcing echo chambers that reduce content diversity, entrench mainstream tastes, and marginalize underrepresented creators and user segments (bakshy2015exposure; wu2024survey). This motivates a paradigm shift toward transparent, distributional user modeling, leading to our central research question:
How can we infer personalized evolving preference distribution via LLM – enhancing interpretability, fairness, and diversity?
To address this problem, we propose Preference Evolution Tracking (PET), a framework that customizes LLMs into probabilistic inference engines for modeling users’ evolving preference distributions. By leveraging pre-softmax logits over a stable preference cluster lattice (e.g., categories or tags), PET captures structured relevance signals more effectively than volatile item-level spaces. This approach enhances interpretability and diversity, especially for long-tail preferences. Specifically, PET supports two scalable inference strategies: Likelihood-based Probing, which iteratively queries the model for each cluster to build a precise distribution, and the more efficient Generative Classification, which extracts all probabilities in a single forward pass. Furthermore, PET employs Hierarchical Probing to maintain tractable inference in extreme multi-label scenarios through semantic taxonomy traversal. Theoretically, the ranking according to the inferred PET distribution guarantees an optimal ranking results (e.g., with respective to NDCG/Recall) under standard assumptions. Applied to domain-aligned LLMs, PET produces portable, interpretable user profiles for long-horizon ranking, fairness auditing, and long-tail preference modeling.
To validate PET’s effectiveness, we conduct comprehensive experiments across three diverse datasets with varying cluster complexity: MovieLens, Yelp, and a large-scale dataset from a world-leading short video platform. Our results demonstrate that the distributional approach significantly outperforms direct-generation baselines and state-of-the-art (SOTA) ranking models on public benchmarks. More critically, on the real-world data of the short video platform, our framework shows a unique ability to rank users’ niche, long-tail interests, significantly outperforming a SOTA model in production. Notably, our goal is not to replace mature production models on highly popular (head) content. Rather, PET complements them by recovering user-specific long-tail preferences that are typically underserved by popularity-optimized systems. Across all experiments, our methods prove more adept at capturing stable, long-term preferences over transient, short-term interests. Our primary contributions are threefold:
  1. 1.
    A New Paradigm for User Modeling: We introduce a novel framework that shifts the focus from optimizing short-term item predictions to generating interpretable, dynamic probability distributions over a stable lattice of human-understandable preference clusters.
  2. 2.
    A Portable and Dynamic User Profile: The learned preference distribution serves as a rich, multi-faceted user profile that captures long-tail interests and is readily portable for diverse downstream tasks, such as user segmentation, trend analysis, or fairness audits.
  3. 3.
    Theoretical and Empirical Validation: We prove that PET’s distributional ranking is theoretically superior to direct generation under standard assumptions. Empirically, we validate PET on three datasets (MovieLens, Yelp, short-video platform), achieving $40\%$ NDCG over direct generation and $20\%$ NDCG over a SOTA ranking model (Qwen3-Reranker-8B (qwen3embedding)) on public benchmarks, and a $7\times$ improvement in long-tail ranking against a SOTA production-level ranking baseline on a real-world dataset.
Figure 1: The PET framework pipeline. Left (Training): An LLM is trained on user history to learn preference distributions. Center (Inference): probing methods extracts the predicted preference distribution from the model’s internal logits. Right (Application): The transparent distribution is used for downstream tasks: personalized ranking, long-tail discovery, and interpretable user profiling.
2 Related Work
User Preference Modeling. To deliver effective personalization, user models aims to capture preferences that are both diverse and dynamic, through three key lines of work. Multi-interest models represent users via multiple embeddings linked to distinct interests (shi2023everyone; cen2020controllable; zhou2018deep). Sequential models, such as FPMC (rendle2010factorizing), GRU4Rec (hidasi2015session), SASRec (kang2018self), and BERT4Rec (sun2019bert4rec), focus on temporal dynamics to predict the next item in a user’s history. More recently, distributional modeling has become central in alignment and fairness, with methods like GDPO for group fairness (yao2024no) and RLHF-based techniques like DPL and AOT for reward uncertainty (siththaranjan2023distributional; melnyk2024distributional). While powerful, these alignment-based methods introduce significant practical considerations, as they often require costly preference-labeled datasets and complex, computationally intensive pipelines.
LLMs in Preference Modeling. Recent work has increasingly applied Large Language Models (LLMs) to the complex task of user preference modeling. Early efforts leveraged LLMs as powerful components for embedding generation (U-BERT; qiu2021u) or unified the field under a prompt-based, text-to-text paradigm (P5; geng2022recommendation). This led to end-to-end generative models like OneRec (deng2025onerec) and (ngo2024recgpt) that directly produce ranked outputs or the most relevant items (wang2025user; wang2024llms), often incorporating sophisticated alignment modules. In a parallel effort toward explainability, RecGPT uses the LLM to generate descriptive, textual user profiles (yi2025recgpt). Crucially, the direct generation of ranked lists via LLM – the approach that forms the basis of our primary baseline – has been shown to consistently outperform traditional sequential models (like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec)) in various settings (geng2022recommendation; wu2024survey; zhao2024recommender; deng2025onerec).
Recent approaches also tackle complex scenarios through structured prompting and multi-context modeling, such as LLM4MSR (wang2024llm4msr), HUM (bao2025heterogeneous), and user interest exploration via latent clusters (wang2024llms). While effective for short-horizon prediction, these methods typically encode a user’s state implicitly within the model’s parameters, without surfacing an interpretable or temporally grounded preference distribution. Our work is also related to logit-probing and generative classification techniques for extracting model beliefs from pre-softmax logits (petroni2019language; schick2020s; raffel2020exploring; zhou2024quest), adapting them to infer evolving user preference distributions from fine-tuned LLMs.
3 Theoretical and Methodological Framework
In this section, we formalize the problem of inferring dynamic, comprehensive user preferences and introduce the Preference Evolution Tracking (PET) framework. We start by outlining the fundamental components of the proposed framework (see Figure 1).
3.1 Problem Formulation
We consider a setting in which the user’s interaction history up to time $t$ is represented as a sequence $H_{1:t}$. The user’s preference at any discrete time $t$ is defined as a probability distribution $\theta_{t}$ over a fixed, finite set of $K$ preference clusters $C:=\{c_{1},\cdots,c_{K}\}$:
$\theta_{t}:=(\theta_{t}(1),\cdots,\theta_{t}(K)),\text{ with }\theta_{t}(i)\geq 0\;\forall i\in[K]\text{ and }\sum_{i\in[K]}\theta_{t}(i)=1.$ (1)
We adopt a latent utility model (mcfadden1972conditional; luce1959individual; train2009discrete), a standard formulation in choice modeling, to describe how this true distribution arises. We assume that there exists a vector of unobserved “attractiveness” scores $\bm{q}:=(q_{1},\cdots,q_{K})$, where $q_{i}$ represents the true strength of the user’s preference for cluster $c_{i}$. The true preference distribution $\theta_{t}$ is then generated via a softmax function:
$\theta_{t}(i)=\frac{\exp(q_{i})}{\sum_{j\in[K]}\exp(q_{j})},\forall i\in[K].$ (2)
Our primary goal is to learn a function, parameterized by an LLM $\mathcal{M}$, that maps this interaction history to a prediction of the user’s future preference distribution, $\hat{\theta}_{t+1:T}$, for a given prediction window $\{t+1:T\}$. We choose to model preferences over this stable cluster space rather than the item space, as it is more robust to the high volatility and enormous scale of items in real-world digital ecosystems.
Since this true distribution $\theta_{t+1:T}$ is not directly observable, we construct an empirical proxy for it from the user’s interactions in a time window, which we denote as $\bar{\theta}_{t+1:T}$. Formally, for each cluster $c_{i}\in C$, its probability is the normalized frequency of interactions:
[…]
To ensure reproducibility and simplicity, we use this transparent and normalized frequency of a single action instead of heuristic weighting of different interaction types, as our framework is designed to be agnostic to the specific construction of the ground-truth label. Formally, we aim to minimize the expected divergence between these two distributions:
$\min_{\mathcal{M}}\mathbb{E}_{H_{1:t}}\left[\mathcal{D}\left(\bar{\theta}_{t+1:T}\parallel\hat{\theta}_{t+1:T}\right)\right],$ (4)
where $\mathcal{D}$ is a measure of distributional similarity, such as the Jensen-Shannon Divergence (JS-Div). The ranked list of preferences used for evaluation with metrics like NDCG is then obtained by sorting the elements of $\hat{\theta}_{t+1:T}$ in descending order.
3.2 Methodology
This section outlines the PET framework for learning user preferences, which consists of two core stages: model alignment and preference inference.
The alignment stage follows a structured two-phase training approach. First, the model undergoes large-scale pre-training (PT) on domain-specific corpora to establish foundational knowledge and capture population-level patterns. Subsequently, it is refined through supervised fine-tuning (SFT) on carefully curated user interaction sequences, effectively aligning its internal representations with the nuances of personalized preference prediction. The SFT data are constructed from each user’s full history by generating pairs (context, label), defined as follows: the context is an expanding historical window $H_{1:t}$, and the label is the user’s future preference distribution $\bar{\theta}_{t+1:T}$. This aligned model forms the basis for all subsequent preference inference methods.
We now present two inference methods for preference distribution estimation, which differ primarily in their querying strategy (more detailed design of the prompts and algorithms presented in Appendices A and B).
Method 1: Likelihood-based Probing. This approach decomposes a multi-class problem into a series of independent binary classifications. The LLM is used as a “prober” or “scorer.” For each cluster $c\in C$, we iteratively ask the LLM for inferences and extract the logit score of positive tokens and then apply a softmax function on the score vector to recover the inferred probability distribution (petroni2019language; schick2020s).
Method 2: Generative Classification. This method uses a single forward pass. A prompt is constructed that frames the task as a multi-choice question, with each cluster mapped to a unique token (e.g., ‘A’, ‘B’). We then read the model’s next-token logits for these specific cluster tokens and normalize them via a softmax function to obtain the final probability distribution (raffel2020exploring; chung2024scaling).
Extreme multi-label preference inference.
To scale PET to real-world applications involving thousands of fine-grained preference clusters – such as in large video platforms or e-commerce systems – we introduce a two-stage Hierarchical Probing strategy. Existing methods face clear limitations in such settings: Likelihood-based Probing exhibits linear computational complexity in the number of clusters $K$, becoming prohibitively expensive, while Generative Classification often yields unstable and semantically ungrounded logits when forced to assign probabilities to artificial class tokens.
To overcome these issues, we organize clusters into a two-level semantic taxonomy: coarse-grained L1 categories (e.g., Food & Drink, Technology) partition the full L2 label space into interpretable subgroups. This tree‑structured approach is also validated in recent extreme multi‑label classification work in other tasks (chen2023token; wan2023explore; zhou2024quest). Inference proceeds in two stages: (1) L1 Preference Scoping: estimates marginal probabilities over top-level categories; (2) Strategic L2 Exploration: selectively probes child clusters, guided by downstream goals. For example, one may prioritize top-ranked L1 branches for relevance, or intentionally explore long-tail branches for novelty. This decomposition improves both tractability and semantic alignment by guiding the LLM from general categories to relevant fine-grained clusters, enabling scalable and adaptive preference modeling. See the detailed computational complexity analysis in Appendix C.
3.3 Theoretical Foundations and Guarantees
This section provides a theoretical argument that PET’s methodology is optimal for standard ranking metrics compared to the direct generation paradigm. We begin with a key assumption about our model:
Assumption 1 (Isotonic Probing Estimator).
We assume that the vector of logit scores $S$ produced by PET’s probing methods (i.e., Algorithms 2 and 1) is order-preserving (isotonic) with respect to the true latent attractiveness scores $\bm{q}$. Formally, for any two clusters $c_{i},c_{j}$, we have $q_{i}\geq q_{j}\Rightarrow\mathbb{E}[S[i]]\geq\mathbb{E}[S[j]].$
1 formalizes the goal that a successfully trained model should learn scores that are correctly ordered according to the true underlying preferences, which is a foundation for many modern preference learning models (guo2017calibration; desai2020calibration; kadavath2022language; schick2020exploiting). With this, we can state the following:
Lemma 1 (Optimality of PET’s Ranking).
With 1, the ranking produced by PET by sorting its inferred probabilities maximizes the expected score for common order-aware metrics, including NDCG@k, Recall@k, and Precision@k. (See proof at Appendix D.)
Remark 1 (Sub-optimality of Direct Generation).
The sub-optimality of direct generation stems from its autoregressive nature, where the ranking is an emergent property of a sequence generation task rather than a direct readout of preference scores. This process is susceptible to multiple known issues: exposure bias, where the model’s training distribution differs from its inference distribution (ranzato2015sequence); calibration errors, where output probabilities are skewed by factors like token frequency (zhao2021calibrate); and search errors from the decoding heuristics themselves, such as beam search (wiseman2016sequence).
4 Experiment
4.1 Experiment Design and Setup
Our experimental design is crafted to comprehensively evaluate the efficacy of PET’s ability in discerning and predicting user preferences across varying granularities and long-term and short-term dynamics. We focus on using two primary inference methodologies for preference prediction and benchmark them against established and SOTA baselines.
Datasets. We evaluate PET on three real-world datasets of varying scale and complexity. MovieLens (harper2015movielens): a widely-used dataset for movie recommendations, representing a relatively small and well-defined genre (cluster) space with 19 clusters (i.e., movie genres). In total, we used 200k users with 33 million rating histories associated with 87k movies. Yelp: a large-scale dataset with 2 million users and 8 million ratings across 1,311 business categories, providing a significantly higher-dimensional preference space. A Real-World Short-Video Dataset: this proprietary dataset is derived from the a real-world short video platform and captures large-scale real-world user interaction behaviors in a highly dynamic environment, reflecting diverse and rich user preference patterns. To ensure privacy, all identifiers are rigorously anonymized. The dataset comprises 50 million implicit feedback records from 16k randomly sampled users, covering 15 million items. Each record contains only anonymized user IDs, item IDs, and interaction types.
[…]
Models and Methods. We use two families of open-source 8B-parameter language models: Qwen3-8B (qwen3technicalreport) and DeepSeek-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability), selected for their strong balance of SOTA performance and computational efficiency. Each model is trained with our PT+SFT (LoRA fine-tune (hu2022lora)) pipeline. This setup supports a comprehensive ablation of training strategies. Our experiments evaluate the primary inference methodologies detailed in Section 3: Likelihood-based Probing (Algorithm 1) and Generative Classification (Algorithm 2). For the Yelp dataset with large cluster space, we specifically evaluate our Hierarchical Probing approach (Algorithm 3). For all methods, the final inferred probability distribution is used to generate a ranked list of clusters for evaluation.
Baselines. We compare PET with two strong LLM-based baselines. To isolate the effect of our inference paradigm, the first baseline Direct Generation uses the same pre-trained and fine-tuned LLM as PET but is prompted to directly generate a top-k ranked list of preferences without producing a probability distribution (geng2022recommendation; deng2025onerec; wang2025user; wang2024llms). As this method does not produce a probability distribution, it is evaluated only on ranking metrics. The second baseline is Qwen3-Reranker-8B, a SOTA pre-trained model specifically optimized for ranking tasks (qwen3embedding). We use it out of the box without any additional fine-tuning. Given a user history and a candidate cluster space, it produces both a ranked list and a probability distribution. For the short-video dataset, we compare against the platform’s production Transformer-based ranker model.
Evaluation Metrics. Model performance is comprehensively evaluated using established metrics compared to long-term (LT) and short-term (ST) ground truths, with a focus on ranking quality and distributional similarity. Ranking metrics include NDCG@k, which assesses the quality of ranking order with an emphasis on top positions; Recall@k, measuring the proportion of relevant items captured within the top k results; and Precision@k, evaluating the exactness of top-k recommendations. Distribution similarity is quantified using Jensen-Shannon Divergence (JS-Div), where lower values indicate closer alignment between predicted and actual preference distributions. All results are averaged across test samples and reported per context window size as specified in the dataset.
Figure 2: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions). Prediction windows: long-term and short-term. Metrics: NDCG@$1,5,10$ and JS-Divergence. We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (PT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model. Note that for the Direct Generation baseline, the JS-divergence will not be available. Sample size: 38,434.
4.2 How Effective and Robust is PET across Different Context Horizon?
We begin by evaluating PET on the MovieLens dataset with 19 clusters, aiming to systematically test its effectiveness and robustness across following key dimensions: (1) different input history lengths and (2) prediction horizons (short-term vs. long-term). We vary the input context window from 1 to 8 sessions, and assess performance on both short-term (next-session) and long-term (multi-session) preference prediction.
As shown in Figure 2, the results demonstrate the robustness and effectiveness of PET across a wide range of settings. The performance gains are especially pronounced in full-list metrics (NDCG@5, NDCG@10), where PET’s distributional inference enables richer modeling of user preferences beyond top-1 accuracy. For example, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 = 0.863, with a +55% improvement over Direct Generation and +20% gain over Qwen3-Reranker-8B. It also improves distributional alignment, reducing JS-Div to 0.316, a 22% relative reduction compared to the Qwen3-Reranker-8B’s 0.406. While Direct Generation is occasionally competitive at NDCG@1, its performance deteriorates as the number of predicted clusters increases, as it might suffers from exposure bias, noted in Remark 1. Our Generative Classification method also shows consistent improvements over Direct Generation on Qwen3-8B base model. Though slightly less accurate than Logit-Probing, it offers a compelling latency–accuracy trade-off due to its single-pass nature, making it a strong candidate for real-time or resource-constrained environments. (See more experimental results in other settings in Appendix F)
To demonstrate PET’s ability to build interpretable profiles, we sampled 125 users and divided their histories into four equal time periods. Using PET’s Likelihood-based Probing, we inferred a distribution over 19 genres for each user in each period. Aggregating these across users yielded group-level preference distributions that evolve over time – visualized in both a heatmap and 3D ribbon plot (Figure 3). These results show that user preference distributions are not static but migrate in nuanced ways. Dominant genres like Drama, Comedy, and Action exhibit stability across periods, representing structural, long-term user interests. In contrast, genres such as Adventure, Romance, and Sci-Fi demonstrate significant temporal fluctuations – highlighting PET’s ability to capture volatile, context-dependent preferences. Notably, genres like Sci-Fi, Fantasy, and IMAX show gradual upward shifts, possibly reflecting emerging or re-surfacing interests. Together, these dynamics validate PET’s unique capability not only in preference prediction, but also in generating descriptive, evolving group-level user profiles – an essential feature for diagnosing user behavior shifts, designing cold-start interventions, or understanding preference stability versus exploration.
4.3 Can PET Handle Thousands of Clusters?
To evaluate PET’s scalability in high-dimensional settings, we use the Yelp dataset with 1,311 fine-grained categories. Since exhaustive Logit-Probing becomes intractable at this scale, we adopt a two-stage Hierarchical Probing strategy (Algorithm 3). First, we construct a semantic hierarchy that maps all categories into 26 high-level L1 clusters (e.g., “Food & Restaurants,” “Health & Medical”). For demonstration, we focus on the high-traffic L1 cluster “Food & Restaurants,” further decomposed into 19 L2 sub-clusters (e.g., “Mexican,” “Vegan”). During inference, PET first identifies a user’s coarse L1 interest, then performs fine-grained ranking within the selected L2 cluster. While operators may target different L1 clusters depending on downstream needs, this experiment provides one representative use case. (See detailed mapping construction in Appendix E.)
Table 1 shows that PET with Hierarchical Probing achieves strong performance across both coarse (L1) and fine-grained (L2) levels. At L1, Logit-Probing is unequivocally dominant. It achieves high and robust NDCG scores (e.g., NDCG@1,5,10 are 0.980, 0.879, and 0.915 for Qwen3-8B), demonstrating an exceptional ability to identify a user’s general interests. In stark contrast, both the Direct Generation and Generative Classification methods fail at this stage, with performance often near zero. These results reinforce the necessity of iterative probing in high-dimensional settings where broad disambiguation is critical. At L2, where the ranking task becomes more localized, both Logit-Probing and Generative Classification significantly outperform the Direct Generation benchmark. Notably, while Logit-Probing remains the top performer for Qwen3-8B (0.666 NDCG@10), Generative Classification emerges as the clear winner for the DeepSeek-distill-Llama-8B model (0.574 NDCG@10), suggesting that decoding strategies for fine-grained inference can be model-dependent.
Together, these findings demonstrate that PET, equipped with hierarchical probing, scales effectively to large cluster spaces. The coarse-to-fine process not only preserves tractability but also allows flexible adaptation across model architectures — with Logit-Probing excelling in breadth and Generative Classification offering a strong, efficient alternative at finer levels of granularity.
(a) Group-level preference evolution as a probability heatmap
(b) Group-level preference evolution as a bar chart
Figure 3: Evolution of group-level movie genre preferences as probability distribution across four time periods. Dataset: Movielens, 125 users over 4 periods. Method: PET Likelihood-based Probing with PT+SFT on Qwen3-8B. Note: due to limited space, we only show top-10 clusters (genres) here.
Qwen3-8B DeepSeek-distill-Llama-8B
Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
Logit Probing .980 .879 .915 .925 .856 .880
Generative Classification .107 .118 .163 .001 .062 .119
L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
Logit Probing .513 .562 .666 .261 .268 .363
Generative Classification .466 .349 .454 .684 .553 .574
Table 1: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task. Results are shown for a context window of 8 using the PT+SFT pipeline. Best NDCG scores for each level are in bold. L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”. Alternative taxonomies can be easily integrated depending on platform objectives. Sample Size: 5000.
4.4 How Does PET Perform on Real-World Short-Video Platform?
We conclude our empirical study with a large-scale, private short-video dataset from a major content platform, comprising 78 content clusters. This experiment is designed to test our method’s efficacy in a real-world environment against a heavily optimized, production-level Transformer-based ranking model. The primary challenge in such an ecosystem is overcoming the inherent feedback loop, where users primarily interact with content they are already shown. Therefore, our evaluation focuses on PET’s ability to complement an existing production model that is heavily optimized for popular, high-frequency head content. In contrast, PET is designed to surface niche, long-tail interests that are often suppressed by popularity bias. We aggregate interactions into daily sessions (a 24-hour window), ensuring a denser and more comprehensive view of a user’s preferences. All results use the PT+SFT pipeline with Qwen3-8B. Following industry practice, we use play duration – the total time a user spends watching each content item – as a robust and continuous proxy for engagement. (See more details in Section E.2)
Results in Table 2 showcase PET’s effectiveness in promoting long-tail interest, diversity, and personalized recommendation. We highlight the setting of 30-day history to 14-day prediction, as it represents the most challenging long-range forecasting task. The production model fails to identify users’ long-tail preferences, with an NDCG@20 of just 0.0243 on the long-tail segment. PET with Logit-Probing achieves NDCG@20 of 0.1971 on the same tail – a +711% improvement. This demonstrates PET’s ability to surface underrepresented, user-specific content that is typically overlooked by popularity-driven production models. Notably, results are consistent for $k\in\{1,5,10,20\}$ with PET achieving more gains compared to the production model when $k$ is small. We also compare across prediction horizons and observe that PET slightly outperforms the baseline more significantly when the forecasting window is longer (14 days vs. 7 days), consistent with our earlier findings at Section 4.2 regarding LLMs’ strength in modeling long-term user preferences. Finally, PET’s performance remains robust across different context lengths. It maintains stable ranking quality even with limited user history, underscoring its ability to effectively summarize user interests with sparse or noisy data.
Taken together, these findings validate PET as a practical and scalable complement to production systems. enhancing long-tail coverage and personal relevance without sacrificing overall performance. Its ability to enhance diversity and personalization – especially for the most engaged users – marks a valuable addition to real-world recommender systems.
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 NDCG@20
14 Days 7 Days SOTA Production (Production) .007 .009 .012 .034
Logit-Probing .113 .124 .147 .203
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .110 .126 .150 .208
30 Days 7 Days SOTA Transformer (Production) .007 .009 .012 .034
Logit-Probing .103 .118 .138 .192
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .108 .122 .143 .197
Table 2: Long-tail preference learning performance on the short-video dataset for high-activity users, comparing PET’s Logit Probing against a SOTA production model across various history (i.e., 14 and 30 days) and prediction windows (i.e., 7 and 14 days). All results are from the PT+SFT pipeline with the Qwen3-8B model.
5 Conclusion
In this paper, we introduce Preference Evolution Tracking (PET), a framework that casts LLM-based personalization as distributional preference mapping: PET infers a user’s cluster-level probability vector via Likelihood-based Probing and Generative Classification, and scales to large label spaces through Hierarchical Probing. Under a mild isotonicity assumption, ranking by PET’s inferred probabilities is optimal for standard order-aware metrics. Empirically, PET is robust across datasets and horizons, outperforming strong generative and reranking baselines and substantially improving coverage of long-tail interests on an industrial-scale corpus. Beyond accuracy, PET yields interpretable, temporally evolving user profiles that enable group-level user analysis and preference-shift detection. Future work includes extending PET to model richer user representations by incorporating auxiliary signals such as demographic attributes, behavioral embeddings, and multimodal content. These enhancements aim to enable more personalized, adaptive, and interpretable preference modeling in real-world recommendation systems.
Appendix
Appendix A Algorithms
Algorithm 1 Likelihood-based Probing.
1:
[…]
9:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
11:for $j\leftarrow 1$ to $K$ do
12:  Use $T_{\text{probe}},H_{1:t},c_{j}$ to construct a prompt $p_{j}$. $\triangleright$ Create prompt for cluster $c_{j}$
13:  $\mathbf{z}_{j}\leftarrow\mathcal{M}(p_{j})$ $\triangleright$ Get logit vector $\mathbf{z}_{j}\in\mathbb{R}^{|V|}$ for prompt $p_{j}$
14:  $s_{+}\leftarrow\sum_{v\in V_{+}}(\mathbf{z}_{j}[v])/|V_{+}|$ $\triangleright$ Calculate mean logit score for all affirmative tokens
15:  $s_{-}\leftarrow\sum_{v\in V_{-}}(\mathbf{z}_{j}[v])/|V_{-}|$ $\triangleright$ Calculate mean logit score for all negative tokens
16:  $S[j]\leftarrow\frac{\exp(s_{+})}{\exp(s_{+})+\exp(s_{-})}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
17:end for
[…]
Algorithm 2 Generative Classification.
1:
[…]
8:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
10:Use $T_{\text{gen}},H_{1:t},C$ to construct a prompt pp.
[…]
12:for $j\leftarrow 1$ to $K$ do
13:  $S[j]\leftarrow\mathbf{z}[v_{j}]$ $\triangleright$ Extract the raw logit for the token of cluster $c_{j}$
14:end for
[…]
Algorithm 3 Hierarchical Probing with Strategic Exploration.
1:
[…]
7:Predicted preferences $\hat{\theta}$ over the L2 clusters $C_{L2}$.
8:
[…]
10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm 1).
[…]
12:
[…]
16:for each selected cluster $c_{j,L1}$ in $C_{L1,\text{selected}}$ do
17:  $C_{L2,\text{subset}}\leftarrow\text{map}(c_{j,L1})$ $\triangleright$ Get L2 children for this L1 parent.
18:  Construct a conditional prompt for $C_{L2,\text{subset}}$ (e.g., “Given interest in $c_{j,L1}$, …”).
19:  Use the conditional prompt to get scores $S_{L2,\text{subset}}$ via “Likelihood-based Probing”.
20:  $P_{L2|L1}\leftarrow\text{softmax}(S_{L2,\text{subset}})$ $\triangleright$ Get conditional probabilities $P(C_{L2}|c_{j,L1})$.
21:  for each cluster $c_{k}$ in $C_{L2,\text{subset}}$ do
22:   $S_{L2}[k]\leftarrow P_{L2|L1}[c_{k}]\times P_{L1}[j]$ $\triangleright$ Apply chain rule for joint probability.
23:  end for
24:end for
25:
[…]
Appendix B Prompt of Inference
Prompt for Likelihood-based Probing Algorithm 1
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, do they like {GENRE} movies? Answer in “Yes” or “No”.
Prompt for Generative Classification Algorithm 2
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, which genre do they like MOST? Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$1$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, tell me the cluster they like the MOST. Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$3$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, rank the top 3 genres they like the most. Answer with the letter only (A, B, C, etc.):
Appendix C Computational Complexity Analysis
Notation.
Let $K$ be the number of preference clusters and $T$ the prompt length (tokens). Let $C(\mathcal{M},T)$ denote the cost of a single forward pass of the language model $\mathcal{M}$ on a prompt of length $T$ (the dominant compute). Each cluster $c$ is associated with a fixed “verbalizer” (token or short phrase) used to read its score from next-token logits.
Complexity of PET Methods
Likelihood-based Probing Algorithm 1.
Our implementation is iterative: we issue a focused prompt per cluster, effectively scoring clusters independently. The total cost scales linearly with the number of probes:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
Generative Classification Algorithm 1.
This method is single-shot: one prompt, one forward pass, then read the next-token logits for the $K$ verbalizers and normalize. The cost is constant with respect to $K$:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)\big).$
Hierarchical Probing Algorithm 3.
This two-stage method targets large $K$. Let $K_{1}$ be the number of coarse L1 clusters, $B$ the number of L1 branches explored, and $K_{2,\text{avg}}$ the average number of L2 sub-clusters per explored branch. Using iterative probing at both levels, the total cost is the sum of L1 scoping and L2 exploration:
$\text{Cost}\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
Since typically $K_{1}\ll K$ and $B\cdot K_{2,\text{avg}}\ll K$, this is substantially cheaper than flat iterative probing $O\!\big(K\cdot C(\mathcal{M},T)\big)$.
Complexity of Baselines (for comparison)
Direct Generation (top-$k$ list).
One prompt with autoregressive decoding of approximately $O(k)$ output tokens:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+k\big).$
If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end latency increases linearly with the number of cutoffs.
Cross-encoder Reranker.
Score each $(\text{history},\text{cluster})$ pair independently:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
Practical Notes
  • KV cache reuse. For iterative methods, reusing the encoded history reduces repeated compute; most marginal cost is in the final token(s) where logits are read.
  • Batching. Batching improves throughput for iterative probing and reranking, but tail latency still scales with the number of batches/probes.
  • Multi-$k$ evaluation. PET’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
Appendix D Proofs
For Lemma 1:
Proof.
By 1, the scores $S$ from PET preserve the order of the true latent attractiveness scores $\bm{q}$. Since the softmax function is monotonic, the inferred probability distribution $\hat{\theta}$ also preserves this order. Thus, sorting by PET’s inferred probabilities produces a ranking, $\pi_{\text{PET}}$, that is identical to the optimal ranking, $\pi^{*}$, which sorts by the true latent scores $\bm{q}$. For a given user and within the time window of $\{t+1:T\}$, we let the vector of attractiveness score $\bm{q}$ be the true latent relevance score of a user for the cluster space. And we denote the sorted (in descending order) vector of $\bm{q}$ as $\tilde{\bm{q}}$ with $\tilde{q}_{1}\geq\cdots\geq\tilde{q}_{K}$. By 1, the retried logit score vector preserves the same order $S$ as $\bm{q}$. Thus, as the softmax function of $S$ preserves the same order of $S$, then by the procedure of PET – sorting the inferred probability distribution in the descending order – the outputted rank by the PET’s procedure has the exactly the same order as the $\tilde{\bm{q}}$.
On the other hand, let’s visit the definition of three ranking metrics. For any given ranking/permutation of cluster set $C$, denoted as $\pi$ (with the associated score vector $q_{\pi}$), we define
  • Recall@k: $\{\text{relevant items in top}-k\}/K$.
[…]
  • NDCG@k: $\text{DCG}(\pi)@k/\text{IDCG}@k$, where $\text{DCG}(\pi)@k:=\sum_{r+1}^{K}\frac{q_{\pi}(r)}{\log_{2}(r+1)}$ and $\text{IDCG}@k$ is the best possible $\text{DCG}(\pi)@k$ over all possible permutation $\pi$.
It is trivial to see that, all these three ranking metrics are maximized when the permutation $\pi$ has the same order as $\tilde{\bm{q}}$. Since the outputted rank by PET has the same order as $\tilde{\bm{q}}$, then PET achieves the optimal ranking scores with respect to these three ranking methods, which gives equal or better rank scores than the direct generation method. ∎
Appendix E Detailed Experimental Design
E.1 Yelp Dataset
On the Yelp dataset, we evaluate PET in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach. In this regime, Logit Probing becomes prohibitively expensive due to the large output space. To address this, we adopt Hierarchical Probing (Algorithm 3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”). For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”. In inference, PET first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category. While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
Hierarchical Probing Mapping. We ask Gemini-2.5 Pro and GPT-5 to produce this mapping with the prompt:
You are an expert in ontology construction, taxonomy design, and user preference modeling, with extensive experience in building interpretable, semantically meaningful cluster hierarchies for preference learning. I will provide you a list of raw user preference clusters, each with an associated weight or frequency representing its importance or user interaction volume (e.g., number of views). Bicycles: number of reviews xxx, average ratings xxx;
Massage: number of reviews xxx, average ratings xxx;
Your task is to: 1. Group these raw clusters into high-level categories (Level-1) that are semantically meaningful and interpretable. 2. Assign each raw cluster to one and only one Level-1 category, forming a two-level hierarchy. 3. Return the result as a JSON object, where the keys are Level-1 category names and the values are lists of Level-2 clusters (i.e., raw cluster names). You should aim for: 1. 20 to 26 Level-1 categories total. 2. Coherent semantics within each group. 3. Human-readable, concise names for each Level-1 category (e.g., “Food & Drink”, “Fitness”, “Beauty”).
With the mapping, of which we verified that Level-1 categories are well-balanced, with no single group dominating the total frequency or collapsing into overly broad labels:
L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, Pets, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
E.2 Real-World Short-Video Dataset
On the short-video dataset, we select the play duration as the proxy metric to represent the user interests. Compared to sparse binary signals like likes or saves, play duration offers a more reliable and verifiable measure of user interest. While some approaches construct weighted combinations of multiple engagement signals, we deliberately opt for a single, interpretable metric to avoid introducing volatility and platform-specific heuristics. The ground-truth distribution is defined as the aggregated play duration across future windows (3, 7, or 14 days). Although we cannot release raw correlation statistics due to data sensitivity and privacy constraints, internal audits confirm strong alignment between play duration and user feedback.
Our analysis targets high-activity users ($\geq$21 active days in the past month), as they represent the platform’s core, most engaged cohort. Since we only have $30$-days of user data, we pick high-activity users for their dense acitivities.
Durations are normalized per user to form a probability distribution over clusters, preventing heavy-usage magnitude from confounding preference shares.
Appendix F Additional Experimental Results
F.1 Evaluation of PET on Movielens
Inference procedure of Direct Generation. Unlike PET, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation .813 .573 .644 .609 .440 .540
B: Logit-Probing .781 .776 .823 .320 .734 .786 .840 .373
C: Gen. Class. .795 .725 .736 .547 .562 .606 .632 .462
Qwen3-Reranker-8B .814 .713 .728 .399
Short A: Direct Generation .701 .506 .571 .499 .369 .450
B: Logit-Probing .667 .675 .740 .466 .542 .619 .714 .514
C: Gen. Class. .638 .661 .789 .538 .499 .490 .501 .551
Qwen3-Reranker-8B .650 .566 .652 .525
3 Long A: Direct Generation .820 .573 .644 .601 .428 .533
B: Logit-Probing .815 .815 .858 .309 .758 .812 .812 .370
C: Gen. Class. .830 .760 .758 .547 .637 .624 .624 .462
Qwen3-Reranker-8B .809 .714 .727 .401
Short A: Direct Generation .702 .495 .554 .492 .361 .453
B: Logit-Probing .686 .674 .707 .460 .501 .612 .712 .507
C: Gen. Class. .661 .631 .671 .539 .492 .499 .507 .545
Qwen3-Reranker-8B .624 .563 .645 .525
5 Long A: Direct Generation .840 .538 .629 .604 .428 .542
B: Logit-Probing .830 .815 .825 .315 .763 .817 .867 .367
C: Gen. Class. .847 .777 .764 .547 .423 .528 .551 .463
Qwen3-Reranker-8B .802 .709 .725 .403
Short A: Direct Generation .723 .496 .546 .493 .359 .454
B: Logit-Probing .716 .720 .781 .460 .496 .606 .708 .507
C: Gen. Class. .673 .635 .676 .535 .489 .498 .507 .546
Qwen3-Reranker-8B .606 .552 .641 .525
8 Long A: Direct Generation .848 .530 .613 .605 .425 .546
B: Logit-Probing .831 .822 .863 .316 .774 .823 .872 .367
C: Gen. Class. .846 .760 .762 .542 .362 .518 .542 .465
Qwen3-Reranker-8B .801 .701 .721 .406
Short A: Direct Generation ..711 .499 .529 .492 .359 .456
B: Logit-Probing .710 .719 .780 .462 .490 .609 .711 .507
C: Gen. Class. .657 .632 .673 .530 .486 .497 .509 .547
Qwen3-Reranker-8B .609 .522 .638 .528
Table 3: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B. Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation .996 .847 .779 .979 .857 .803
B: Logit-Probing .993 .951 .754 .993 .966 .919
C: Gen. Class. .992 .890 .736 .948 .821 .760
Short A: Direct Generation .921 .549 .484 .772 .467 .417
B: Logit-Probing .887 .711 .589 .788 .688 .608
C: Gen. Class. .882 .627 .517 .772 .588 .434
3 Long A: Direct Generation .999 .857 .773 .972 .845 .799
B: Logit-Probing .991 .960 .914 .996 .973 .926
C: Gen. Class. .997 .911 .792 .897 .854 .755
Short A: Direct Generation .926 .555 .488 .775 .466 .429
B: Logit-Probing .909 .738 .617 .755 .693 .618
C: Gen. Class. .902 .654 .518 .775 .601 .446
5 Long A: Direct Generation .998 .866 .767 .976 .842 .804
B: Logit-Probing .996 .961 .913 .993 .973 .931
C: Gen. Class. .998 .906 .789 .869 .857 .756
Short A: Direct Generation .930 .559 .488 .768 .463 .429
B: Logit-Probing .926 .749 .622 .746 .687 .620
C: Gen. Class. .912 .662 .525 .762 .597 .443
8 Long A: Direct Generation .999 .869 .776 .975 .837 .807
B: Logit-Probing .996 .963 .918 .998 .977 .934
C: Gen. Class. .999 .907 .794 .858 .861 .757
Short A: Direct Generation .925 .555 .486 .774 .461 .431
B: Logit-Probing .922 .749 .621 .736 .694 .622
C: Gen. Class. .907 .658 .521 .762 .597 .447
Table 4: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation .072 .299 .547 .071 .301 .563
B: Logit-Probing .072 .342 .637 .073 .350 .658
C: Gen. Class. .072 .317 .552 .068 .289 .533
Short A: Direct Generation .124 .349 .603 .097 .285 .519
B: Logit-Probing .120 .461 .743 .106 .446 .772
C: Gen. Class. .118 .402 .654 .097 .367 .538
3 Long A: Direct Generation .073 .304 .544 .070 .297 .562
B: Logit-Probing .072 .345 .651 .073 .354 .666
C: Gen. Class. .072 .325 .559 .064 .302 .531
Short A: Direct Generation .124 .354 .608 .096 .278 .527
B: Logit-Probing .122 .477 .777 .101 .443 .773
C: Gen. Class. .121 .417 .653 .096 .369 .543
5 Long A: Direct Generation .073 .310 .544 .070 .295 .563
B: Logit-Probing .073 .349 .656 .072 .351 .666
C: Gen. Class. .073 .326 .562 .061 .302 .529
Short A: Direct Generation .125 .357 .609 .095 .278 .529
B: Logit-Probing .125 .486 .784 .099 .441 .780
C: Gen. Class. .123 .423 .662 .095 .368 .542
8 Long A: Direct Generation .073 .309 .547 .070 .292 .564
B: Logit-Probing .073 .348 .655 .073 .353 .668
C: Gen. Class. .073 .324 .562 .060 .304 .529
Short A: Direct Generation .125 .356 .607 .096 .277 .532
B: Logit-Probing .126 .488 .785 .099 .445 .781
C: Gen. Class. .123 .422 .659 .095 .370 .547
Table 5: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold.
F.2 Evaluation of PET on Yelp
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation 0.0015 0.1069 0.2430 0.0127 0.3109 0.3425
B: Logit-Probing 0.8604 0.8436 0.8669 0.6145 0.9581 0.8501 0.8777 0.6499
C: Gen. Class. 0.0945 0.1147 0.1587 0.6969 0.0000 0.1021 0.1437 0.7245
Short A: Direct Generation 0.0149 0.1133 0.2686 0.0127 0.2971 0.3333
B: Logit-Probing 0.7430 0.7467 0.7891 0.6768 0.7072 0.7352 0.7650 0.6969
C: Gen. Class. 0.0398 0.1122 0.1549 0.7301 0.0001 0.0928 0.1468 0.7296
3 Long A: Direct Generation 0.0040 0.1164 0.3381 0.0170 0.3142 0.3450
B: Logit-Probing 0.9557 0.8724 0.9016 0.5976 0.9590 0.8683 0.8866 0.6527
C: Gen. Class. 0.1254 0.1230 0.1684 0.6935 0.0000 0.0647 0.1451 0.7179
Short A: Direct Generation 0.0277 0.1403 0.3556 0.0219 0.3034 0.3379
B: Logit-Probing 0.8743 0.7970 0.8548 0.6611 0.3391 0.4507 0.5194 0.7020
C: Gen. Class. 0.0454 0.1120 0.1577 0.7235 0.0028 0.0552 0.1823 0.7334
5 Long A: Direct Generation 0.0115 0.1279 0.3522 0.0176 0.3136 0.3481
B: Logit-Probing 0.9629 0.8713 0.9075 0.5848 0.7814 0.8057 0.8350 0.6579
C: Gen. Class. 0.1204 0.1243 0.1720 0.6839 0.0004 0.0567 0.1554 0.7143
Short A: Direct Generation 0.1101 0.1472 0.3557 0.0203 0.3001 0.3394
B: Logit-Probing 0.8823 0.7923 0.8585 0.6511 0.6402 0.7237 0.7643 0.6944
C: Gen. Class. 0.0411 0.1137 0.1556 0.7194 0.0905 0.0802 0.1107 0.7317
8 Long A: Direct Generation 0.0081 0.1390 0.3297 0.0156 0.3084 0.3429
B: Logit-Probing 0.9795 0.8787 0.9152 0.5734 0.9246 0.8563 0.8796 0.6546
C: Gen. Class. 0.1070 0.1183 0.1628 0.7012 0.0013 0.0615 0.1187 0.7168
Short A: Direct Generation 0.0739 0.1347 0.3152 0.0176 0.2948 0.3367
B: Logit-Probing 0.8593 0.7874 0.8599 0.6470 0.6809 0.7250 0.7637 0.6984
C: Gen. Class. 0.0402 0.1227 0.1526 0.7341 0.1095 0.0863 0.1292 0.7328
Table 6: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation 0.0054 0.4100 0.2583 0.1423 0.2503 0.2667
B: Logit-Probing 0.9538 0.6749 0.4897 0.9892 0.5025 0.3898
C: Gen. Class. 0.6152 0.3920 0.3456 0.0018 0.3508 0.2656
Short A: Direct Generation 0.0408 0.2252 0.1707 0.0444 0.2130 0.1885
B: Logit-Probing 0.8613 0.3718 0.2754 0.7905 0.3910 0.2628
C: Gen. Class. 0.1453 0.2006 0.1947 0.0006 0.1976 0.1681
3 Long A: Direct Generation 0.0186 0.3880 0.2929 0.1531 0.2516 0.2672
B: Logit-Probing 0.9886 0.6379 0.4957 0.9946 0.5557 0.4017
C: Gen. Class. 0.7713 0.3863 0.3482 0.0006 0.2281 0.2606
Short A: Direct Generation 0.0828 0.2101 0.2060 0.0576 0.2172 0.1902
B: Logit-Probing 0.9526 0.3905 0.3186 0.4106 0.2403 0.2018
C: Gen. Class. 0.1477 0.1963 0.1983 0144 0.1005 0.1634
5 Long A: Direct Generation 0.0690 0.3539 0.2918 0.1609 0.2549 0.2881
B: Logit-Probing 0.9898 0.6293 0.5030 0.9040 0.5651 0.4384
C: Gen. Class. 0.7863 0.4104 0.3592 0.0018 0.1878 0.1997
Short A: Direct Generation 0.3854 0.1894 0.2015 0.0540 0.2172 0.1975
B: Logit-Probing 0.9616 0.3864 0.3268 0.7419 0.4715 0.3124
C: Gen. Class. 0.1339 0.2110 0.1962 0.3025 0.1028 0.0850
8 Long A: Direct Generation 0.0755 0.3220 0.3002 0.1519 0.2533 0.2757
B: Logit-Probing 0.9970 0.6293 0.5070 0.9676 0.5442 0.4164
C: Gen. Class. 0.6783 0.3903 0.3338 0.0078 0.2077 0.2478
Short A: Direct Generation 0.2636 0.1887 0.2048 0.0462 0.2108 0.1954
B: Logit-Probing 0.9652 0.3844 0.3413 0.8025 0.4415 0.2920
C: Gen. Class. 0.1429 0.2394 0.1838 0.3938 0.1122 0.1185
Table 7: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation 0.0007 0.2977 0.3807 0.0173 0.1994 0.4070
B: Logit-Probing 0.1614 0.5224 0.7448 0.1675 0.4038 0.6063
C: Gen. Class. 0.0881 0.2823 0.5051 0.0002 0.2653 0.3910
Short A: Direct Generation 0.0089 0.2678 0.4435 0.0099 0.3070 0.4955
B: Logit-Probing 0.2604 0.5066 0.7148 0.2319 0.5153 0.6654
C: Gen. Class. 0.0304 0.2305 0.4533 0.0002 0.2313 0.3975
3 Long A: Direct Generation 0.0025 0.2850 0.4488 0.0190 0.2032 0.4138
B: Logit-Probing 0.1699 0.4997 0.7660 0.1706 0.4500 0.6296
C: Gen. Class. 0.1149 0.2773 0.5120 0.0001 0.1672 0.3889
Short A: Direct Generation 0.0180 0.2496 0.5500 0.0117 0.3146 0.5038
B: Logit-Probing 0.2899 0.5252 0.8171 0.1288 0.3325 0.5308
C: Gen. Class. 0.0304 0.2214 0.4585 0.0030 0.1171 0.4046
5 Long A: Direct Generation 0.0096 0.2624 0.4490 0.0196 0.2012 0.4378
B: Logit-Probing 0.1666 0.4903 0.7741 0.1532 0.4516 0.6763
C: Gen. Class. 0.1181 0.2950 0.5330 0.0002 0.1340 0.2853
Short A: Direct Generation 0.0926 0.2283 0.5344 0.0113 0.3070 0.5138
B: Logit-Probing 0.2859 0.5151 0.8345 0.2182 0.6206 0.7940
C: Gen. Class. 0.0272 0.2392 0.4568 0.0700 0.1174 0.1942
8 Long A: Direct Generation 0.0105 0.2411 0.4689 0.0183 0.2012 0.4208
B: Logit-Probing 0.1729 0.4966 0.7887 0.1645 0.4377 0.6477
C: Gen. Class. 0.1016 0.2811 0.4921 0.0013 0.1484 0.3565
Short A: Direct Generation 0.0605 0.2223 0.5327 0.0097 0.3088 0.5182
B: Logit-Probing 0.2870 0.5137 0.8638 0.2392 0.5892 0.7530
C: Gen. Class. 0.0301 0.2784 0.4256 0.0908 0.1279 0.2726
Table 8: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold. Level: L1