### v3

# SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference

Abstract
Large Language Models (LLMs) are increasingly used to understand user preferences, typically via the direct generation of ranked item lists. However, this end-to-end generative paradigm inherits the bias and opacity of autoregressive decoding, over-emphasizing frequent (head) preferences and obscure long-tail ones, thereby biasing personalization toward head preferences. To address this, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), which treats the LLM as an implicit probabilistic model by probing it to infer a probability distribution over interpretable preference clusters. In doing so, SPECTRA reframes user modeling from sequence generation with decoding heuristics to distributional inference, yielding explicit, cluster-level user preference representations. We evaluate SPECTRA on MovieLens, Yelp, and a large-scale short-video platform, demonstrating significant gains across three dimensions: SPECTRA achieves (i) distributional alignment, reducing Jensen–Shannon divergence to empirical distributions by $25\%$ against strong baselines; (ii) long-tail exposure, reducing decoding-induced head concentration and increasing global exposure entropy by $30\%$; and (iii) downstream applications such as personalized ranking, translating these gains into a $40\%$ NDCG boost on public datasets and a 7× improvement on ranking long-tail preferences against an industry-leading Transformer-based production baseline.
SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference
Luyang Zhang thanks: Part of this work was done while at TikTok Inc. Carnegie Mellon University luyangz@andrew.cmu.edu          Jialu Wang TikTok Inc. faldict@ucsc.edu          Shichao Zhu TikTok Inc. shichao.szhu@gmail.com
Beibei Li Carnegie Mellon University beibeili@andrew.cmu.edu          Zhongcun Wang TikTok Inc.          Guangmou Pan TikTok Inc.
Yang Song TikTok Inc. ys@sonyis.me
1 Introduction
Understanding user preferences in an explicit and interpretable manner is a central challenge in building personalized AI systems (Doshi-Velez and Kim, 2017; Amershi et al., 2019; Rudin, 2019; Lipton, 2018). Large Language Models (LLMs) have emerged as a promising tool due to their ability to leverage rich semantic signals from textual side information (e.g., reviews, descriptions) alongside interaction logs. However, moving beyond simple pattern matching benefits from treating the LLM not merely as an autoregressive generator of item candidates or ranked lists from textualized user histories, but as an implicit probabilistic model over latent user intent, in line with emerging theoretical frameworks that interpret in-context learning as approximate Bayesian inference over latent variables (Bengio et al., 2025; Wang et al., 2023; Li et al., 2023). This motivates seeking explicit, time-varying preference distributions over interpretable semantic clusters that can be analyzed, audited, and reused across downstream applications.
The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked lists of items (Wang et al., 2025, 2024a; Ngo and Nguyen, 2024; Yu et al., 2024; Deng et al., 2025). Even recent methods adopting textual rationales or chain-of-thought ultimately rely on this autoregressive decoding process to produce final recommendations (Wang et al., 2024b; Yi et al., 2025). While effective for predicting short-term interactions with already-popular item, this paradigm leaves critical gaps. First, it keeps the latent preference representation implicit: generation fails to expose an explicit, structured posterior over preferences, making it infeasible for audits and explanations. Second, in direct generation, reliance on decoding heuristics (e.g., beam search, top-$k$ sampling) could exacerbate head concentration by truncating low-probability clusters. As detailed in Section˜3.3, this structural vulnerability to frequency bias and search truncation might suppresses the long tail (Holtzman et al., 2020; Stahlberg and Byrne, 2019), preventing faithful personalization of user-specific niche preferences.
These limitations have significant consequences for user-facing AI systems. The lack of an explicit user model hampers transparency and governance, making it difficult to systematically inspect or audit how preferences are represented across users and groups. At the same time, systematically suppressing low-probability but relevant items reinforces popularity during training and deployment, reducing content diversity and marginalizing long-tail preferences on real-world platforms (Bakshy et al., 2015; Wu et al., 2024). Together, these observations motivate a shift from sequence generation to explicit, distributional user modeling, leading to our central research question:
How can we infer personalized preference distributions via LLMs as explicit posteriors over semantic clusters that preserve diverse user preferences and support downstream decision-making?
To address this question, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), a distributional user modeling framework that treats the LLM as an implicit probabilistic model and extracts a cluster-level probability vector over a semantically interpretable set of preference clusters by probing the model’s logits. Using a cluster vocabulary (e.g., tags or categories) provides a stable semantic interface, avoiding the volatility and sparsity of item-level identifiers while yielding preference representations that are comparable across users and time. By representing preferences as distributions – rather than decoded item lists – SPECTRA reduces head-dominated concentration and better preserves niche preferences, improving coverage and exposure diversity in the resulting preference representation. For scalability to large cluster space, SPECTRA remains tractable via coarse-to-fine inference over a semantic taxonomy.
To validate SPECTRA, we evaluate on three datasets spanning varying scale and cluster complexity: MovieLens, Yelp, and a large-scale industrial short-video platform. Across these settings, SPECTRA consistently outperforms direct generation (DG) and strong LLM-based ranking baselines, improving distributional alignment, exposure diversity, and downstream ranking quality. Beyond short-horizon next-step prediction, our experiments further show that SPECTRA more reliably captures persistent, long-term preference structure over time, yielding stronger performance in long-horizon evaluation settings. SPECTRA achieves particularly significant gains on under-represented preference segments, consistent with its ability to mitigate head-dominated concentration; accordingly, SPECTRA is not intended to replace mature head-optimized production rankers, but to complement them by recovering user-specific niche preferences that are often suppressed under decoding-based pipelines. Our primary contributions are threefold:
  1. 1.
    Distributional user modeling from LLMs. We reframe LLM-based user modeling from autoregressive list generation to distributional inference, producing cluster-level preference distributions over a semantically interpretable cluster vocabulary.
  2. 2.
    Scalable posterior extraction via probing. We extract these cluster-level distributions by probing LLM logits, and design inference procedures that remain tractable for large cluster vocabularies via coarse-to-fine hierarchical inference over a semantic taxonomy.
  3. 3.
    Extensive Empirical Validation. Across MovieLens, Yelp, and a large-scale industrial short-video platform, SPECTRA yields consistent gains in distributional fidelity and coverage: it improves alignment to empirical preference distributions (e.g., $23\%$ JS divergence vs. a strong LLM-based ranking baseline such as Qwen3-Reranker-8B (QwenTeam, 2025)) and increases exposure diversity (up to $30\%$ global exposure entropy) relative to direct generation. As a downstream consequence, these representation gains also improve ranking outcomes (up to $40\%$ NDCG on public benchmarks) and deliver up to 7× higher NDCG on under-represented preference segments than a production Transformer baseline.
Figure 1: The SPECTRA framework pipeline. Left (Training): we adapt an LLM (CPT + SFT) to predict a user’s future preference distribution over $K$ clusters from historical context. Center (Inference): we extract cluster scores from the model’s logits via likelihood-based probing or generative classification (yielding score vector $S$, and obtain a predicted distribution $\hat{\theta}=\text{softmax}(S)$ (and ranking via $\text{argsort}(S)$). Application: the transparent distribution supports personalized ranking, long-tail interest discovery, and interpretable preference profiling.
2 Related Work
User Preference Representation. Prior work models user preferences from interaction histories using multi-preference representations (Shi et al., 2023; Cen et al., 2020; Zhou et al., 2018) and sequential next-item predictors (Rendle et al., 2010; Hidasi et al., 2015; Kang and McAuley, 2018; Sun et al., 2019). These methods typically encode user state implicitly (as embeddings or item-level scores), which limits transparency for analysis and auditing (Doshi-Velez and Kim, 2017; Lipton, 2018; Rudin, 2019). An open challenge is to surface an explicit, cluster-level preference distribution that is directly comparable across time.
LLMs for Preference Representation. LLMs have been incorporated into preference representation via prompt-based unification (P5; (Geng et al., 2022)) and text-aware representation learning (U-BERT; (Qiu et al., 2021)), enabling direct list generation and generative ranking (Ngo and Nguyen, 2024; Deng et al., 2025; Wang et al., 2024a, 2025) as well as textual profile generation for explainability (Yi et al., 2025). Recent work enriches this with review-grounded preference cues (Kim et al., 2025), user-embedding contextualization (Ning et al., 2025), and LLM-enhanced generative retrieval (Paischer et al., 2025). Reasoning-style pipelines (e.g., CoT4Rec, LLMRG) generate item-level rationales for recommendations (Yue et al., 2025; Wang et al., 2024b). These approaches typically keep user state implicit (in hidden representations or text) or operate at the item level. In contrast, SPECTRA addresses the upstream task of extracting a time-comparable, cluster-level preference distribution that can serve as a structured signal for downstream rankers or reasoning modules.
Belief Extraction and Scalable Inference. Belief extraction from token likelihoods has been studied mainly for factual knowledge probing and prompt-based classification (Petroni et al., 2019; Schick and Schütze, 2020; Raffel et al., 2020; Zhou et al., 2024). This paradigm can also be applied to infer preference distributions over semantic clusters in user modeling settings. Direct decoding can suffer from decoding/search artifacts (Holtzman et al., 2020; Stahlberg and Byrne, 2019), motivating inference procedures that avoid reliance on decoding heuristics. Distributional objectives in alignment and uncertainty modeling (Siththaranjan et al., 2023; Yao et al., 2024; Melnyk et al., 2024) further motivate predicting an explicit preference distribution. Scaling to large cluster spaces commonly uses coarse-to-fine hierarchical inference, as in extreme multi-label classification (Zhou et al., 2024).
3 Theoretical and Methodological Framework
In this section, we formalize distributional user preference modeling in an interpretable cluster space, describe SPECTRA’s two-stage pipeline (alignment and score-based inference), and provide a brief structural analysis contrasting distributional readout with direct generation (see Figure˜1).
3.1 Task and Setup
Let a user’s context up to time $t$ consist of (i) an interaction history $H_{1:t}$ (e.g., rated items, reviewed businesses, watched videos), (ii) user-side metadata $U$ (e.g., bio/self-intro), and (iii) item-side metadata associated with the interacted items, denoted by $\Phi_{1:t}$ (e.g., categories/tags, prices, creator attributes). We therefore denote the full input context as $X_{1:t}:=\big(H_{1:t},\,U,\,\Phi_{1:t}\big).$
We represent user preferences in a fixed, interpretable cluster space $C=\{c_{1},\dots,c_{K}\}$, where each cluster corresponds to a semantic preference category (e.g., a genre, cuisine, or topic). We model preferences over clusters rather than items because item spaces are large and volatile, whereas cluster semantics are stable and interpretable for tracking and auditing. Given context $X_{1:t}$, our goal is to predict the user’s future preference distribution over $C$ for a prediction window [t+1,T][t\!+\!1,T], denoted by $\theta_{t+1:T}\in\Delta^{K-1}$, and output $\hat{\theta}_{t+1:T}=\hat{\theta}(X_{1:t})$. We interpret $\theta_{t+1:T}$ as a softmax over latent utilities for clusters (equivalently, the multinomial logit form widely used in preference modeling) (McFadden, 1972; Luce and others, 1959; Train, 2009). Since the true $\theta$ is not directly observable, we construct an empirical target $\bar{\theta}_{t+1:T}$ by normalizing interaction counts in the future window: $\bar{\theta}_{t+1:T}(i):=n_{i}/\sum_{j}n_{j}$, where $n_{i}$ is the interaction count for cluster $c_{i}$.
3.2 Methodology
SPECTRA formulates user modeling as distributional inference over an interpretable cluster space. It consists of (1) Alignment, which trains an LLM to map user context to a preference distribution over clusters, and (2) Inference, which reads out this distribution via score-based operators, with a hierarchical strategy for large label spaces. (more detailed design of the methods are presented in Appendices˜A and B)
Alignment (Continuous PT + LoRA SFT).
We adapt the base LLM to the platform domain via continuous pre-training (CPT) on textualized domain data from user metadata, item metadata, and interaction traces $(H_{1:t},U,\Phi_{1:t})$ (e.g., serialized profiles, item titles/tags/descriptions, and interaction text such as reviews or captions). We then apply supervised fine-tuning (SFT) with LoRA on interaction sequences, where the target is a textual encoding of the future empirical preference distribution over clusters (computed from interactions in [t+1,T][t\!+\!1,T]), rather than a single next-item label. This aligns the model to predict multi-interest user preference distributions.
Preference inference as score-based querying.
Given context $X_{1:t}$ and a textual label/definition for each cluster $c_{i}$, SPECTRA estimates a preference distribution over the cluster space $C$ by querying the aligned LLM with pre-specified prompt templates. Concretely, SPECTRA constructs a score vector $S(X_{1:t})\in\mathbb{R}^{K}$ and outputs $\hat{\theta}_{t+1:T}=\mathrm{softmax}(S(X_{1:t}))$. We compute $S$ in two ways:
(1) Likelihood-based probing. For each cluster $c_{i}$, we pose a yes/no probe (Petroni et al., 2019) under the template prompt $\pi_{\text{probe}}(X_{1:t},c_{i})$ and extract the next-token logits $\mathbf{z}_{i}\in\mathbb{R}^{|V|}$. Using small affirmative/negative token sets $V_{+}$ and $V_{-}$, we compute the mean logits $\mu_{i,+}:=\frac{1}{|V_{+}|}\sum_{v\in V_{+}}\mathbf{z}_{i}[v]$ and $\mu_{i,-}:=\frac{1}{|V_{-}|}\sum_{v\in V_{-}}\mathbf{z}_{i}[v]$, and set
$S_{i}:=\mu_{i,+}-\mu_{i,-}.$ (1)
(2) Generative classification. We construct a single multi-choice prompt (Raffel et al., 2020; Chung et al., 2024) $\pi_{\text{gen}}(X_{1:t},C)$ that lists clusters and maps each cluster $c_{i}$ to an answer token $v_{i}$ (e.g., A, B, …). The LLM produces a next-token logit vector $\mathbf{z}\in\mathbb{R}^{|V|}$ for this prompt, and we take $S_{i}:=\mathbf{z}[v_{i}]$.
For both approaches, we eventually normalize the score vector $S$ with a softmax to obtain $\hat{\theta}_{t+1:T}$.
Scaling via hierarchical probing.
For extreme multi-label settings, flat likelihood probing incurs $O(K)$ scoring cost. We introduce hierarchical probing by organizing $C$ into a two-level taxonomy (coarse L1 categories $\rightarrow$ fine-grained L2 clusters). Inference proceeds top-down: (1) L1 scoping: estimate a distribution over L1 categories; (2) L2 exploration: selectively probe only L2 clusters within a subset of L1 branches (e.g., high-mass branches for relevance, or long-tail branches for exploration, depending on the application goal). The explored branch set $\mathcal{G}$ can be chosen to trade off relevance and exploration. This reduces the scoring cost to $O\!\left(K_{\mathrm{L1}}+\sum_{g\in\mathcal{G}}K_{\mathrm{L2}}(g)\right)$ (L1 categories plus the probed L2 clusters under explored branches $g$); details are in Appendix˜C.
Hierarchical probing supports both (i) an automatic top-down expansion under high-probability parent clusters and (ii) optional analyst-guided drill-down to inspect specific preference facets; the latter is used for interpretability rather than improving ranking accuracy.
3.3 Structural Analysis: Distributional Readout vs. Direct Generation
We explain why SPECTRA’s distributional readout is robust relative to direct generation (DG) and summarize the key theoretical implications (proofs in Appendix˜D).
In the idealized limit where the learned cluster scores $S(X_{1:t})$ are perfectly calibrated under cross-entropy, sorting clusters by these scores recovers the Bayes-optimal ordering for common order-aware metrics (e.g., NDCG, Recall) (Gneiting and Raftery, 2007; Goodfellow et al., 2016; Blasiok et al., 2023; Tewari and Bartlett, 2007). In a more realistic regime where the SPECTRA scores are only $\varepsilon$-approximately isotonic (i.e., any preferred pair is misordered with probability at most $\varepsilon$), we show that the expected number of pairwise misorderings is $O(K^{2}\varepsilon)$ (see Appendix˜D).
Why DG degrades in practice. In practice, SPECTRA exhibits structural advantages over decoding-based DG. While both methods use the same underlying LLM, DG is susceptible to three well-documented limitations:
  • Search / truncation. Decoding/search heuristics (e.g., beam search) can exhibit pathological behaviors and impose a hard truncation: valid long-tail clusters that fall outside the explored set are effectively pruned (Stahlberg and Byrne, 2019; Wiseman and Rush, 2016). SPECTRA instead scores clusters directly and normalizes globally, preserving tail mass.
  • Frequency bias. Autoregressive generation can amplify high-frequency patterns from pre-training, biasing outputs toward head concepts (Zhao et al., 2021; Holtzman et al., 2020). SPECTRA normalizes scores over the given cluster set, reducing corpus-level frequency effects.
  • Exposure bias. DG is sequential: early deviations change the context for later predictions, compounding errors (Ranzato et al., 2015; Bengio et al., 2015). SPECTRA extracts cluster scores marginally from the same user context, avoiding error compounding across positions.
Together with the bounds in Appendix˜D, these structural considerations align with our empirical findings: as $K$ and the target list length grow, DG degrades more sharply on long-tail preference prediction, while SPECTRA remains robust.
4 Experiment
4.1 Experimental Design and Setup
We evaluate SPECTRA as a cluster-level user preference modeling method: given a user history, the model infers a probability distribution over interpretable preference clusters, from which a ranked list is derived for downstream evaluation. We report both short-term prediction (next-step preference) and long-term prediction (future preference profile), and evaluate: (i) distributional alignment, (ii) mitigating popularity concentration collapse, and (iii) downstream ranking performance.
Datasets. We use three datasets with varying cluster-space size. MovieLens (Harper and Konstan, 2015) offers a compact, low-dimensional space (K=19K{=}19 genres); we use 200k users with 33M ratings over 87k movies. Yelp represents a high-dimensional semantic space (2M users and 8M ratings across K=1,311K{=}1{,}311 business categories). Finally, we use a large-scale Short-Video dataset (50M interactions from 16k users over 15M items) to evaluate performance in a high-velocity, industrial setting.
Training and evaluation protocol. For MovieLens and Yelp, we order each user’s interactions chronologically and split by time: the first 80% is used for CPT+SFT, and the remaining 20% is held out for evaluation. We report two prediction targets: long-term (LT) preference prediction, where the ground truth is the empirical cluster distribution computed over the held-out future interactions, and short-term (ST) prediction, where the target is the cluster of the next interaction.  [^1] We define a session as a fixed-length block of consecutive interactions: 5 interactions for MovieLens, 10 for Yelp, and one day of interactions (daily aggregation) for the short-video dataset. For the short-video dataset, we use a fixed-window protocol: a 30-day interaction window for CPT+SFT input and the subsequent 14-day period for evaluation.
Models and inference methods. We study two 8B open LMs: Qwen3-8B (QwenTeam, 2025) and DeepSeek-Distill-Llama-8B (DeepSeek-AI, 2025). We fine-tune each base model with our CPT+SFT pipeline using LoRA (Hu et al., 2022). We evaluate our primary inference methods from Section˜3.2: Likelihood-based Probing (Algorithm˜1) and Generative Classification (Algorithm˜2). For Yelp (large $K$), we additionally evaluate Hierarchical Probing (Algorithm˜3). All methods output a cluster distribution, from which we derive a ranked list over the full cluster set.
Baselines. We compare against two strong LLM-based baselines (and a production personalization model for short-video). Direct Generation (DG) uses the same fine-tuned LLM and directly decodes a top-$k$ list of clusters (without inferring an explicit probability distribution) (Geng et al., 2022; Deng et al., 2025; Wang et al., 2025, 2024a); it is therefore evaluated on ranking and diversity metrics.
Qwen3-Reranker-8B (Zhang et al., 2025a) is a pretrained reranker used without additional fine-tuning; it outputs per-cluster relevance scores on the same inputs, which we use for ranking and normalize into a distribution when needed. For the short-video dataset, we additionally compare against the platform’s production Transformer-based personalization model (a behavior-sequence model used for ranking)(Liu et al., 2022; Zhang et al., 2025b; Guan et al., 2025).
Figure 2: MovieLens (Qwen3-8B, CPT+SFT): trends over context length (1/3/5/8 sessions) for long-term (LT; future preference profile) and short-term (ST; next-step preference) prediction. Left: NDCG@10 ($\uparrow$). Right: JS-Div to the empirical LT distribution ($\downarrow$). Direct Generation does not produce a distribution and is evaluated on NDCG only. DeepSeek-Distill-Llama-8B yields similar trends at Appendix˜F.
Evaluation metrics. We report (i) distributional alignment via Jensen-Shannon divergence (JS-Div) between predicted and empirical cluster distributions, (ii) exposure diversity (mitigating popularity concentration) via Global Exposure Entropy  (Adomavicius and Kwon, 2011) computed from the cluster exposure distribution induced by users’ top-$k$ lists, and (iii) ranking performance via NDCG@k, Recall@k, and Precision@k. Full metric definitions and implementation details are provided in Section˜E.1.
Computational Efficiency. Naive Likelihood-based Probing scales linearly with the number of clusters $K$. With KV-cache prefix reuse, we encode the shared user-history prefix once ($T_{\mathrm{pre}}$ tokens) and only recompute a short cluster-specific suffix per probe ($T_{\mathrm{suf}}$ tokens), which substantially reduces overhead when $T_{\mathrm{pre}}\!\gg\!T_{\mathrm{suf}}$ (see Appendix˜C). Our Generative Classification variant is single-shot (independent of $K$), with the same order of compute as a single forward pass.
4.2 Robust Preference Modeling Across Context Lengths and Time Scales
We begin with MovieLens (K=19K{=}19) to test SPECTRA across (i) input context length and (ii) prediction time scale. We vary the context window from 1 to 8 sessions and evaluate both short-term (next-step) and long-term (future-profile) preference prediction.
Distributional alignment. Figure˜2 shows that SPECTRA yields better alignment to empirical preference distributions across context lengths, with particularly clear gains in the long-term setting. (DG does not output a distribution and is therefore not evaluated on JS-Div). For example, under CPT+SFT Qwen3-8B with context window =8=8, Logit-Probing achieves JS-Div =0.316=0.316 versus $0.406$ for Qwen3-Reranker-8B (22% relative reduction), indicating more calibrated cluster-level preference profiles.
Pred. Method Entropy@1 Entropy@3
LT DG 1.020 2.446
Logit-Probing 1.747 3.130
Gen. Class. 1.575 2.732
ST DG 0.810 2.163
Logit-Probing 2.093 3.246
Gen. Class. 1.605 2.696
Table 1: Global Exposure Entropy (higher is better) on MovieLens with CPT+SFT Qwen3-8B, reported at $k\in\{1,3\}$ for long-term (LT) and short-term (ST) prediction. Full results are in Table˜6.
Exposure diversity and mode collapse. We next examine exposure diversity via Global Exposure Entropy (Table˜1). Across both long-term and short-term settings, the DG baseline shows pronounced concentration at small cutoffs (e.g., Entropy@1 =1.020=1.020 under Qwen3-8B, LT), suggesting decoding collapses onto a few globally popular genres. In contrast, SPECTRA maintains substantially higher entropy already at k=1k{=}1 (e.g., $1.747$ in LT and $2.093$ in ST), surfacing user-specific interests at the very top of the ranking. As expected, entropy saturates for larger $k$ due to the finite genre space, making small-$k$ behavior the most diagnostic for concentration effects.
Downstream ranking performance. These profile-level improvements translate into stronger ranking, especially at larger cutoffs. For instance, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 =0.863=0.863, improving over DG (+55%) and Qwen3-Reranker-8B (+20%). While DG can be competitive at top-1, its performance degrades as $k$ increases, consistent with the concentration effects observed above. Our Generative Classification variant also consistently improves over DG while offering a favorable latency-accuracy trade-off via single-pass inference. See detailed metrics including Recall@k, and Precision@k at Appendix˜F.
Figure 3: Group-level genre preference distributions across four time periods (MovieLens; 125 users). Method: SPECTRA Likelihood-based Probing with CPT+SFT Qwen3-8B. Darker cells indicate higher probability mass; only the top-10 genres are shown for space.
Qualitative analysis: inspecting distributional preference shifts. Finally, to illustrate interpretability beyond aggregate metrics, we sample 125 users and split their histories into four equal time periods. Applying SPECTRA’s Likelihood-based Probing yields per-period genre distributions, whose aggregation reveals systematic changes over time (Figure˜5). Dominant genres such as Drama, Comedy, and Action remain relatively stable across periods, while others (e.g., Adventure, Romance, Sci-Fi) fluctuate more substantially. We also observe gradual increases for genres such as Sci-Fi and IMAX in the aggregated profiles, illustrating SPECTRA’s ability to surface nuanced temporal shifts in an explicit, inspectable form.
4.3 Scalable Preference Modeling in High-Dimensional Cluster Spaces
We next evaluate whether SPECTRA scales to high-dimensional cluster spaces using Yelp (K=1,311K{=}1{,}311 categories). Because flat inference over all categories is impractical at this scale, we adopt a coarse-to-fine Hierarchical Probing strategy (Algorithm˜3). We induce a two-level semantic taxonomy by grouping categories into $K_{1}{=}26$ L1 domains (e.g., “Food & Restaurants”, “Health & Medical”) using category text (names/descriptions); SPECTRA first infers an L1 preference distribution, then ranks L2 sub-clusters within a selected L1 branch (mapping in Appendix˜E). As a representative case, we report results on the high-traffic L1 domain “Food & Restaurants,” which contains $K_{2}{=}19$ L2 sub-clusters (e.g., “Mexican”, “Vegan”); other branches can be evaluated analogously. Notably, the taxonomy is used primarily for candidate pruning; because scoring relies on natural-language descriptors, the method the method is robust to arbitrary label IDs and to moderate changes in taxonomy design.
Coarse (L1) identification. Table˜2 shows that hierarchical Logit-Probing is highly effective at identifying users’ coarse interests. For Qwen3-8B, it achieves NDCG@1/5/10 of 0.980/0.879/0.915 at L1, while Direct Generation degenerates (0.008/0.139/0.330), effectively failing to identify the correct coarse domain. This motivates coarse-to-fine inference for large $K$, and shows that the L1 branch selection step benefits most from per-cluster probing.
Level Method NDCG
@1 @5 @10
L1 (26 clusters) DG .008 .139 .330
Logit-Probing .980 .879 .915
Gen. Class. .107 .118 .163
L2 (19 sub-clusters) DG .298 .368 .509
Logit-Probing .513 .562 .666
Gen. Class. .466 .349 .454
Table 2: Hierarchical probing on Yelp (LT, window=8=8) with Qwen3-8B. NDCG at L1 (26 super-clusters) and L2 (19 sub-clusters under “Food & Restaurants”). Best per level in bold; n=5,000n{=}5{,}000. DeepSeek-Distill-Llama-8B shows the same trend; see full table in Section˜F.2.
Fine-grained (L2) ranking within a branch. Conditioned on the selected L1 domain, SPECTRA also improves fine-grained preference ranking among the 19 L2 sub-clusters. At L2, both Logit-Probing and Generative Classification outperform DG on both fine-tuned LLMs; Logit-Probing is strongest for Qwen3-8B (e.g., NDCG@10=0.666=0.666). DeepSeek-Distill-Llama-8B shows the same overall pattern – SPECTRA variants substantially outperform DG at both levels (see Appendix˜F).
4.4 Long-Tail Preference Discovery in Production Short-Video Personalization
We conclude with a large-scale, short-video dataset, comprising 78 platform-defined content clusters and a strong production Transformer-based behavior-sequence personalization model (e.g., Liu et al., 2022; Zhang et al., 2025b; Guan et al., 2025). This setting evaluates SPECTRA in an interaction-dense personalized system where observed engagement is shaped by an exposure feedback loop and production systems are heavily optimized for frequent head interests.We therefore focus on whether SPECTRA can complement the production stack by elevating user-specific niche interests typically under-exposed by the production stack.
Evaluation setup. We aggregate user activity into daily sessions. Given a history window (14 or 30 days), we evaluate predictions on a held-out future window (7 or 14 days). Since user preference is not directly observable under strong recommendation feedback, we use cumulative play duration as a behavioral proxy for engagement and construct the ground-truth cluster profile by summing play duration over items within each cluster in the future window and normalizing across clusters (details in Section˜E.3). SPECTRA outputs a cluster distribution; the production system operates at the item level. We compare them by aggregating the clusters of served items into an exposure distribution (normalized over clusters) for the prediction window.
Long-tail definition. We define the long-tail as clusters in the bottom quartile of historical engagement mass (computed on the training window), i.e., low-prevalence clusters most susceptible to under-exposure (last $25\%$ of the cluster sets, equivalent to 20 of 78 clusters in our data). We report NDCG@k computed specifically on this tail subset, where higher values indicate better recovery of user-specific niche interests.
Context Prediction Method NDCG
@1 @3 @5 @10
14 Days 7 Days Production .007 .009 .012 .034
Logit-Probing .113 .124 .147 .203
14 Days Production .004 .006 .008 .024
Logit-Probing .110 .126 .150 .208
30 Days 7 Days Production .007 .009 .012 .034
Logit-Probing .103 .118 .138 .192
14 Days Production .004 .006 .008 .024
Logit-Probing .108 .122 .143 .197
Table 3: Short-video long-tail performance (high-activity users). We compare SPECTRA (Logit-Probing, Qwen3-8B with PT+SFT) against the platform production model under varying history windows (14/30 days) and prediction windows (7/14 days). Metric: NDCG@{1,3,5,10} on the long-tail cluster subset.
[…]
5 Conclusion
We presented Preference Evolution Tracking (SPECTRA), a framework for distributional user modeling with LLMs. Given a user’s history, SPECTRA infers an explicit posterior over interpretable preference clusters via Likelihood-based Probing or Generative Classification, and scales to large cluster spaces with Hierarchical Probing. In theory, SPECTRA decouples preference inference from sequence decoding by ranking clusters using explicit probability readouts, which avoids decoding-induced exposure bias that can degrade Direct Generation at larger cutoffs. Across public benchmarks and an industrial short-video dataset, SPECTRA is robust across context lengths and time scales, consistently improving distributional alignment, mitigating popularity concentration, and strengthening cluster-level ranking compared to strong baselines. Beyond predictive gains, SPECTRA produces inspectable, temporally evolving preference profiles that support auditing and group-level analysis of preference shifts. Future work includes extending the framework to incorporate additional signals (e.g., multimodal content and auxiliary behavioral representations) and studying robustness under evolving taxonomies and feedback loops in deployed personalization pipelines.
Appendix
Appendix A Algorithms
Algorithm 1 Likelihood-based Probing.
1:
[…]
9:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
11:for $j\leftarrow 1$ to $K$ do
12:  Use $T_{\text{probe}},H_{1:t},c_{j}$ to construct a prompt $p_{j}$. $\triangleright$ Create prompt for cluster $c_{j}$
13:  $\mathbf{z}_{j}\leftarrow\mathcal{M}(p_{j})$ $\triangleright$ Get logit vector $\mathbf{z}_{j}\in\mathbb{R}^{|V|}$ for prompt $p_{j}$
14:  $s_{+}\leftarrow\sum_{v\in V_{+}}(\mathbf{z}_{j}[v])/|V_{+}|$ $\triangleright$ Calculate mean logit score for all affirmative tokens
15:  $s_{-}\leftarrow\sum_{v\in V_{-}}(\mathbf{z}_{j}[v])/|V_{-}|$ $\triangleright$ Calculate mean logit score for all negative tokens
16:  $S[j]\leftarrow s_{+}-s_{-}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
17:end for
[…]
Algorithm 2 Generative Classification.
1:
[…]
8:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
10:Use $T_{\text{gen}},H_{1:t},C$ to construct a prompt pp.
[…]
12:for $j\leftarrow 1$ to $K$ do
13:  $S[j]\leftarrow\mathbf{z}[v_{j}]$ $\triangleright$ Extract the raw logit for the token of cluster $c_{j}$
14:end for
[…]
Algorithm 3 Hierarchical Probing with Strategic Exploration.
1:
[…]
7:Predicted preferences $\hat{\theta}_{t+1:T}$ over the L2 clusters $C_{L2}$.
8:
[…]
10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm˜1).
[…]
12:
[…]
16:for each selected cluster $c_{j,L1}$ in $C_{L1,\text{selected}}$ do
17:  $C_{L2,\text{subset}}\leftarrow\text{map}(c_{j,L1})$ $\triangleright$ Get L2 children for this L1 parent.
18:  Construct a conditional prompt for $C_{L2,\text{subset}}$ (e.g., “Given interest in $c_{j,L1}$, …”).
19:  Use the conditional prompt to get scores $S_{L2,\text{subset}}$ via “Likelihood-based Probing”.
20:  $P_{L2|L1}\leftarrow\text{softmax}(S_{L2,\text{subset}})$ $\triangleright$ Get conditional probabilities $P(C_{L2}|c_{j,L1})$.
21:  for each cluster $c_{k}$ in $C_{L2,\text{subset}}$ do
22:   $S_{L2}[k]\leftarrow P_{L2|L1}[c_{k}]\times P_{L1}[j]$ $\triangleright$ Apply chain rule for joint probability.
23:  end for
24:end for
25:
[…]
Appendix B Prompt of Inference
Prompt for Likelihood-based Probing Algorithm˜1
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, do they like {GENRE} movies? Answer in “Yes” or “No”.
Prompt for Generative Classification Algorithm˜2
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, which genre do they like MOST? Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$1$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, tell me the cluster they like the MOST. Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$3$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, rank the top 3 genres they like the most. Answer with the letter only (A, B, C, etc.):
Appendix C Computational Complexity Analysis
Notation.
Let $K$ be the number of preference clusters and $T$ the total prompt length (tokens). Let $C(\mathcal{M},T)$ denote the cost of a single forward evaluation of the language model $\mathcal{M}$ on $T$ tokens (dominant compute). For KV/prefix reuse, we decompose each probing prompt into a shared prefix of length $T_{\mathrm{pre}}$ (user history + fixed scaffold) and a cluster-specific suffix of length $T_{\mathrm{suf}}$, so that $T=T_{\mathrm{pre}}+T_{\mathrm{suf}}$. We write $C(\mathcal{M},T_{\mathrm{pre}})$ and $C(\mathcal{M},T_{\mathrm{suf}})$ for the corresponding costs.
Verbalizers. In generative classification, each cluster $c$ is associated with a verbalizer (e.g., a single token/letter) whose next-token logit is read and normalized. In likelihood probing, the cluster is inserted into the prompt and we read the next-token logits of a fixed response verbalizer set (e.g., Yes/No).
Complexity of SPECTRA Methods
Likelihood-based Probing Algorithm˜1.
Our probing implementation is iterative: we score clusters via $K$ cluster-conditioned prompts. Without reuse, this requires $K$ full prompt evaluations:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
With prefix/KV reuse, we evaluate the shared prefix once and only re-evaluate a short cluster-specific suffix per probe:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T_{\mathrm{pre}})+K\cdot C(\mathcal{M},T_{\mathrm{suf}})\big).$
[…]
Generative Classification Algorithm˜2.
This method is single-shot: one prompt, one forward pass, then read the next-token logits for the $K$ verbalizers and normalize. The cost is constant with respect to $K$:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)\big).$
Hierarchical Probing Algorithm˜3.
This two-stage method targets large $K$. Let $K_{1}$ be the number of coarse L1 clusters, $B$ the number of L1 branches explored, and $K_{2,\text{avg}}$ the average number of L2 sub-clusters per explored branch. Using iterative probing at both levels, the total cost is the sum of L1 scoping and L2 exploration:
Cost $\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)$
$\quad\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
Since typically $K_{1}\ll K$ and $B\cdot K_{2,\text{avg}}\ll K$, this is substantially cheaper than flat iterative probing $O\!\big(K\cdot C(\mathcal{M},T)\big)$.
Complexity of Baselines (for comparison)
Direct Generation (top-$k$ list).
One prompt with an initial prefill plus autoregressive decoding of $m(k)$ output tokens (typically $m(k)=O(k)$ under a constrained letter-list format):
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M})\big),$
where $C_{\mathrm{dec}}(\mathcal{M})$ denotes the per-token decode cost given a cached prefix (implementation-dependent). If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end cost scales linearly with the number of cutoffs.
Cross-encoder Reranker.
Score each $(\text{history},\text{cluster})$ pair independently:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
Practical overhead (token-length proxy).
Define the relative compute overhead of probing (with reuse) versus a single-shot method as
$\rho\;=\;\frac{C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}})}{C(\mathcal{M},T_{\mathrm{pre}}+T_{\mathrm{suf}})}.$
In typical prompting setups, $T_{\mathrm{pre}}$ (user history) can dominate and $T_{\mathrm{suf}}$ is relatively short, so prefix reuse can reduce the overhead from $\approx K\times$ to a constant factor that depends on the ratio $T_{\mathrm{pre}}/T_{\mathrm{suf}}$. Absolute latency depends on batching and serving implementations; the proxy above is intended to summarize scaling rather than provide measured runtimes.
Method # model evals No reuse With prefix/KV reuse
Direct Generation (top-$k$) 11 prefill + $m(k)$ decode $\approx O(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M}))$ (same)
Generative Classification 11 $O(C(\mathcal{M},T))$ (same)
Likelihood Probing $K$ $O(K\,C(\mathcal{M},T))$ $O(C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}}))$
Table 4: Complexity summary. $m(k)$ is the number of generated tokens in direct generation.
Example regimes (token-length proxy).
For illustration, consider K=20K{=}20 and compare probing with reuse ($T_{\mathrm{pre}}+KT_{\mathrm{suf}}$) to a single-shot pass ($T_{\mathrm{pre}}+T_{\mathrm{suf}}$) under two regimes: (i) short history: $T_{\mathrm{pre}}{=}80$, $T_{\mathrm{suf}}{=}20$, giving a ratio $\frac{80+20\cdot 20}{80+20}=\frac{480}{100}\approx 4.8\times$; (ii) long history: $T_{\mathrm{pre}}{=}300$, $T_{\mathrm{suf}}{=}15$, giving $\frac{300+20\cdot 15}{300+15}=\frac{600}{315}\approx 1.9\times$. These token-length ratios are proxies for scaling; actual wall-clock depends on batching and kernels.
Regime $T_{\mathrm{pre}}$ $T_{\mathrm{suf}}$ $K$ Proxy ratio $\frac{T_{\mathrm{pre}}+KT_{\mathrm{suf}}}{T_{\mathrm{pre}}+T_{\mathrm{suf}}}$
[…]
Table 5: Token-length proxy ratios for probing with prefix/KV reuse (illustrative; not measured latency).
Practical Notes
  • KV cache reuse. For iterative methods, reusing the encoded history reduces repeated compute; most marginal cost is in the final token(s) where logits are read.
  • Batching. Batching improves throughput for iterative probing and reranking, but tail latency still scales with the number of batches/probes.
  • Multi-$k$ evaluation. SPECTRA’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
Appendix D Extra Theories and Proofs
For the theoretical analysis, we consider an idealized preference mapping $g_{\phi}:H_{1:t}\rightarrow\mathbb{R}^{K}$ that produces logits $\boldsymbol{S}(H_{1:t})=g_{\phi}(H_{1:t})$ and an induced distribution $\hat{\theta}_{t+1:T}(H_{1:t})=\mathrm{softmax}(\boldsymbol{S}(H_{1:t}))$. We assume that $\phi$ is chosen to minimize the population cross-entropy between the latent-softmax model and the predictor:
$\mathcal{L}_{\mathrm{CE}}(\phi)=\mathbb{E}_{H_{1:t}}\big[\mathrm{CE}\big(\theta_{t+1:T},\hat{\theta}_{t+1:T}\big)\big].$ (2)
Training on empirical labels $\bar{\theta}_{t+1:T}$ can be viewed as a finite-sample approximation to this population objective. For the theoretical analysis in this appendix, we work in an idealized population setting. We posit a latent preference distribution $\theta(H_{1:t})=\mathrm{softmax}(\mathbf{q}(H_{1:t}))$ and view $\bar{\theta}_{t+1:T}$ as a finite-sample Monte Carlo estimate of $\theta(H_{1:t})$ (i.e., $\bar{\theta}_{t+1:T}\approx\theta(H_{1:t})$ as the number of interactions grows). Consequently, the empirical objective with $\bar{\theta}$ approximates the population objective with $\theta$.
D.1 Preliminaries: Idealized Bayes-optimality
We first establish the behavior of the model in an idealized population limit. Let $\boldsymbol{S}\in\mathbb{R}^{K}$ be the model’s logits, produced by SPECTRA’s probing methods (i.e., Algorithms˜2 and 1). The model predicts $\hat{\theta}=\mathrm{softmax}(\mathbf{S})$, and we train by minimizing cross-entropy between $\theta$ and $\hat{\theta}$. This follows the standard view of cross-entropy (log-loss) as a strictly proper scoring rule: at the population optimum, the predictor’s output distribution coincides with the true conditional distribution (Gneiting and Raftery, 2007; Goodfellow et al., 2016; Blasiok et al., 2023). In multiclass settings, this implies Bayes-consistency of the induced classifier (Tewari and Bartlett, 2007).
Proposition 1 (Isotonicity at convergence).
Since cross-entropy is a strictly proper scoring rule (Gneiting and Raftery, 2007; Blasiok et al., 2023), the global minimum of the population risk is achieved if and only if $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$. Equivalently, in the latent-softmax model $\theta=\mathrm{softmax}(\boldsymbol{q})$, the Bayes-optimal predictor recovers the true conditional distribution (Tewari and Bartlett, 2007; Goodfellow et al., 2016). By injectivity of softmax up to an additive constant, this implies $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some scalar $c$. Thus, in the idealized limit, the model’s logits are perfectly order-preserving (isotonic) with respect to latent preferences.
Proof.
Let the true latent preference distribution be $\theta=\mathrm{softmax}(\boldsymbol{q})$ and the predicted distribution be $\hat{\theta}=\mathrm{softmax}(\boldsymbol{S})$. The expected population cross-entropy risk is defined as:
$\mathcal{R}=\mathbb{E}_{H}\left[-\sum_{k=1}^{K}\theta_{k}\log\hat{\theta}_{k}\right].$ (3)
By adding and subtracting the entropy of the true distribution $H(\theta)=-\sum\theta_{k}\log\theta_{k}$, we can rewrite the risk as:
[…]
where $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence. A fundamental property of proper scoring rules (Gibbs’ inequality) states that $D_{\mathrm{KL}}(\theta\|\hat{\theta})\geq 0$, with equality holding if and only if $\hat{\theta}=\theta$ almost everywhere. Therefore, the global minimum is achieved uniquely when $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$.
The softmax function is invariant to constant shifts, meaning $\mathrm{softmax}(\boldsymbol{x})=\mathrm{softmax}(\boldsymbol{y})$ implies $\boldsymbol{x}=\boldsymbol{y}+c\boldsymbol{1}$ for some scalar $c\in\mathbb{R}$. Consequently, at the global minimum, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$.
Finally, since translation by a scalar $c$ does not alter the relative ordering of elements, for any pair of clusters $i,j$:
$S_{i}>S_{j}\iff(q_{i}+c)>(q_{j}+c)\iff q_{i}>q_{j}.$ (5)
This confirms that in the idealized limit, the logits $\boldsymbol{S}$ are perfectly isotonic with the latent utility $\boldsymbol{q}$. ∎
In the language-modeling context, this corresponds to the standard interpretation of cross-entropy training as maximum-likelihood estimation of the next-token distribution and an upper bound on the true entropy rate (Braverman et al., 2020). Our analysis applies this Bayes-optimal perspective to the cluster-distribution predictor used by SPECTRA.
D.2 Robustness: The $\varepsilon$-Approximate Regime
Proposition proposition˜1 characterizes an ideal Bayes-optimal limit where $\boldsymbol{S}$ is exactly isotonic with $\boldsymbol{q}$. In practice, models are only approximately calibrated (Blasiok et al., 2023; Braverman et al., 2020). We capture this via a probabilistic pairwise condition.
Assumption 1 ($\varepsilon$-Approximate Pairwise Isotonicity).
For a given user and time window, let $\boldsymbol{q}\in\mathbb{R}^{K}$ be latent scores and $\boldsymbol{S}\in\mathbb{R}^{K}$ the logits produced by SPECTRA. We say $\boldsymbol{S}$ is $\varepsilon$-approximately pairwise isotonic w.r.t. $\boldsymbol{q}$ if, for all $i,j$ with $q_{i}>q_{j}$,
$\mathbb{P}\Big(S_{i}<S_{j}\Big)\;\leq\;\varepsilon,$ (6)
where the probability is over randomness in training, histories, and probing. That is, each strictly preferred pair is misordered with probability at most $\varepsilon$.
We measure ranking quality by the number of misordered pairs relative to the Bayes-optimal ranking.
Definition 1 (Pairwise ranking regret).
Let $\pi^{*}$ be the permutation that sorts clusters by decreasing $q_{i}$, and let $\pi$ be any permutation of $\{1,\dots,K\}$. Define
$\mathcal{R}(\pi;\boldsymbol{q})=\sum_{\begin{subarray}{c}i,j\in[K]\\ q_{i}>q_{j}\end{subarray}}\mathbb{I}\big(\pi^{-1}(i)>\pi^{-1}(j)\big),$ (7)
where $\pi^{-1}(i)$ denotes the position (rank) of cluster $i$ under the permutation $\pi$ (smaller is better). Thus, $\mathcal{R}(\pi;\boldsymbol{q})$ counts the number of strictly preferred pairs $(i,j)$ that are reversed under $\pi$.
Let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$ in descending order. The following bound is immediate.
Theorem 2 (Approximate optimality of SPECTRA).
Under Assumption 1, the expected pairwise regret of $\pi_{\mathrm{SPECTRA}}$ satisfies
$\mathbb{E}\big[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})\big]\;\leq\;\binom{K}{2}\,\varepsilon=O(K^{2}\varepsilon),$ (8)
where the expectation is over the randomness in $\boldsymbol{S}$.
Proof.
For $q_{i}>q_{j}$, define $E_{ij}=\{\pi_{\mathrm{SPECTRA}}^{-1}(i)>\pi_{\mathrm{SPECTRA}}^{-1}(j)\}=\{S_{i}<S_{j}\}.$ By Assumption 1, $\mathbb{P}(E_{ij})\leq\varepsilon$. By Definition 1, $\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})=\sum_{q_{i}>q_{j}}\mathbb{I}(E_{ij}).$ Taking expectations and using linearity,
$\mathbb{E}[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})]=\sum_{q_{i}>q_{j}}\mathbb{P}(E_{ij})\leq\sum_{q_{i}>q_{j}}\varepsilon\leq\binom{K}{2}\varepsilon.$ (9)
Thus, if SPECTRA is “mostly isotonic” in the sense of Assumption 1 (small $\varepsilon$), its ranking is close to Bayes-optimal: only $O(K^{2}\varepsilon)$ pairs are misordered in expectation. In the limit $\varepsilon\to 0$, we recover the exact isotonicity of Proposition 1.
D.3 From cross-entropy to the $\varepsilon$ regime
The previous result treats $\varepsilon$ as a behavioral property of the SPECTRA head. We now connect $\varepsilon$ to the training objective $\mathcal{L}_{\mathrm{CE}}(\phi)$.
Let $\mathcal{L}_{\mathrm{CE}}^{\star}=\mathcal{L}_{\mathrm{CE}}(\theta_{t+1:T})$ denote the Bayes-optimal risk, achieved when $\hat{\theta}_{t+1:T}=\theta_{t+1:T}$, and define the excess cross-entropy risk
$\Delta_{\mathrm{CE}}$ $=\mathcal{L}_{\mathrm{CE}}(\phi)-\mathcal{L}_{\mathrm{CE}}^{\star}$ (10)
$=\mathbb{E}_{H_{1:t}}\big[\mathrm{KL}\big(\theta_{t+1:T}(H_{1:t})\,\|\,\hat{\theta}_{t+1:T}(H_{1:t})\big)\big].$ (11)
By Pinsker’s inequality and Jensen’s inequality, we obtain:
Lemma 1 (L1 error from excess cross-entropy).
The expected $L_{1}$ distance between the latent and predicted preference distributions is bounded by
$\mathbb{E}_{H_{1:t}}\big[\|\theta_{t+1:T}(H_{1:t})-\hat{\theta}_{t+1:T}(H_{1:t})\|_{1}\big]\;\leq\;$
$\quad\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (12)
Proof.
Recall that the excess cross-entropy risk is exactly the expected Kullback-Leibler divergence between the true distribution $\theta$ and the predicted distribution $\hat{\theta}_{t+1:T}$:
[…]
For any specific history $H_{1:t}$, Pinsker’s inequality bounds the $L_{1}$ distance by the KL divergence:
$\|\theta-\hat{\theta}\|_{1}\;\leq\;\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}$. (14)
Taking the expectation over histories $H_{1:t}$ on both sides:
$\mathbb{E}\left[\|\theta-\hat{\theta}\|_{1}\right]\;\leq\;\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]. (15)
Since the square root function $f(x)=\sqrt{x}$ is concave, Jensen’s inequality implies that $\mathbb{E}[\sqrt{X}]\leq\sqrt{\mathbb{E}[X]}$. Applying this to the right-hand side:
$\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]\; $\leq\;\sqrt{2\,\mathbb{E}\left[D_{\mathrm{KL}}(\theta\|\hat{\theta})\right]}\;$ (16)
$=\;\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (17)
Combining these steps yields the lemma. ∎
As $\Delta_{\mathrm{CE}}\to 0$ (e.g., with more data, capacity, and optimization), the predicted distributions $\hat{\theta}_{t+1:T}$ converge to $\theta_{t+1:T}$ in expected $L_{1}$ distance. Intuitively, if $\hat{\theta}_{t+1:T}$ is close to $\theta_{t+1:T}$, then it is unlikely to significantly distort the ordering between clusters with a clear preference gap. Formalizing $\varepsilon$ as an explicit function of $\Delta_{\mathrm{CE}}$ for all pairs would require additional global margin assumptions on $\theta_{t+1:T}$, which are often unrealistic in long-tail regimes. Instead, we interpret $\varepsilon$ as capturing residual misorderings after training: Lemma˜1 shows that cross-entropy minimization drives $\hat{\theta}_{t+1:T}$ toward $\theta_{t+1:T}$, thereby pushing the model into a small-$\varepsilon$ regime for most histories and clearly preferred pairs.
D.3  SPECTRA vs. Decoding-Based Direct Generation
We now compare SPECTRA to decoding-based direct generation (DG), which uses the same logits $\boldsymbol{S}$ but produces a ranked list via an autoregressive decoding algorithm (e.g., greedy, beam, sampling).
Let $\pi^{*}\in\Pi_{K}$ be the permutation that sorts clusters by decreasing $\boldsymbol{q}$, and let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$. We model any decoding-based ranking as follows.
Definition 2 (Decoding-based ranking).
A decoding-based ranking procedure is a (possibly randomized) algorithm that maps logits $\boldsymbol{S}\in\mathbb{R}^{K}$ to a permutation of clusters. We denote by
$\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$
the (random) permutation produced by such a decoder when run on $\boldsymbol{S}$. The randomness here comes from any sampling or tie-breaking used by the decoding algorithm (e.g., top-$p$ sampling, temperature, stochastic beam search). We write $\mathbb{E}_{\mathrm{DG}}[\cdot]$ for expectation with respect to this decoder randomness, conditioned on the logits $\boldsymbol{S}$.
We measure quality using a generic order-aware loss.
Assumption 2 (Order-aware ranking loss).
Let $\mathcal{L}:\Pi_{K}\times\mathbb{R}^{K}\to\mathbb{R}$ be a loss such that, for any fixed $\boldsymbol{q}$:
  1. 1.
    $\mathcal{L}(\pi^{*};\boldsymbol{q})\leq\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$;
  2. 2.
    for almost all $\boldsymbol{q}$ (w.r.t. any continuous distribution), the minimizer is unique: $\mathcal{L}(\pi^{*};\boldsymbol{q})<\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\neq\pi^{*}$.
In the Bayes-optimal regime of Proposition 1, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, SPECTRA recovers $\pi^{*}$ by simple sorting. We show that no decoding procedure can do better on top of these logits.
Theorem 3 (Dominance of SPECTRA over decoding-based generation).
Fix $\boldsymbol{q}\in\mathbb{R}^{K}$ and assume $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some $c$. Let $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}$ and let $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$ be the (possibly random) permutation produced by any decoding-based procedure when run on logits $\boldsymbol{S}$. Under Assumption 2,
$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (18)
where the expectation is over the decoder’s internal randomness. Moreover, if $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$ and the decoder outputs any $\pi\neq\pi^{*}$ with nonzero probability, then the inequality is strict.
Proof.
Since $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, sorting $\boldsymbol{S}$ is equivalent to sorting $\boldsymbol{q}$: $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}=\mathrm{argsort}\boldsymbol{q}=\pi^{*}.$ By Assumption 2, $L(\pi^{*};\boldsymbol{q})\leq L(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$.
For any fixed realization of the decoder’s randomness, the decoder outputs some permutation $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$, and thus
$\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\;\geq\;\mathcal{L}(\pi^{*};\boldsymbol{q})=\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}).$ (19)
Taking expectation over the decoder’s randomness gives
$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (20)
since $\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})$ is deterministic.
If $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$, then whenever $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})=\mathcal{L}(\pi^{*};\boldsymbol{q})$ we must have $\pi_{\mathrm{DG}}(\boldsymbol{S})=\pi^{*}$. Thus equality in expectation can hold only if the decoder returns $\pi^{*}$ almost surely. If instead the decoder outputs some $\pi\neq\pi^{*}$ with nonzero probability, then on that event $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})>\mathcal{L}(\pi^{*};\boldsymbol{q})$, which makes the inequality strict. ∎
Theorem 3 is an idealized statement: when logits are perfectly calibrated ($\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$), SPECTRA’s simple logit-sorting is Bayes-optimal, and any decoding-based mapping on top of the same logits can at best match, but not improve, the induced ranking. In practice, SPECTRA operates in the $\varepsilon$-approximate regime of Assumption 1, where its regret is controlled by Theorem 2, while decoding-based methods incur additional errors from exposure bias, truncation, and search heuristics.
Remark 1 (Risk Decomposition: SPECTRA vs. Direct Generation).
Theorem 3 characterizes an idealized regime where logits are perfectly calibrated ($\boldsymbol{S}\approx\boldsymbol{q}$). In the practical regime, Direct Generation (DG) incurs additional error sources that SPECTRA avoids due to its distributional nature. We formally conceptualize the ranking risk of DG as:
$\mathrm{Risk}(\text{DG})$ $\;\approx\;\mathrm{Risk}(\text{SPECTRA})\;+\;\Delta_{\text{freq}}$
$\quad\;+\;\Delta_{\text{trunc}}\;+\;\Delta_{\text{exp}}.$ (21)
Here, $\mathrm{Risk}(\text{SPECTRA})$ is controlled by the approximate isotonicity parameter $\varepsilon$ (Theorem 2), while the additive penalty terms represent the structural limitations identified in Section 3.3:
  1. 1.
    $\Delta_{\text{freq}}$ (Frequency Bias): Represents the divergence caused by the autoregressive model’s tendency to over-produce high-frequency tokens from pre-training (Zhao et al., 2021; Holtzman et al., 2020), which SPECTRA mitigates via closed-set normalization.
  2. 2.
    $\Delta_{\text{trunc}}$ (Search/Truncation Error): Represents the expected utility loss from heuristic search (e.g., beam search), which imposes a hard truncation on valid long-tail clusters falling outside the beam (Wiseman and Rush, 2016; Stahlberg and Byrne, 2019).
  3. 3.
    $\Delta_{\text{exp}}$ (Exposure Bias): Represents the error accumulated due to context distribution shift during autoregressive decoding (Ranzato et al., 2015).
This decomposition theoretically grounds our empirical observation that SPECTRA is significantly more robust in long-tail ranking tasks.
Appendix E Detailed Experimental Design
E.1 Setups
Global Exposure Entropy. Let $\mathcal{U}$ be the set of evaluation users and $\mathcal{C}$ the cluster space (e.g., genres). For a given method and cutoff $k$, let $R_{u,k}\subseteq\mathcal{C}$ denote the set of clusters appearing in the top-$k$ list for user $u\in\mathcal{U}$. We define the global exposure count for cluster $c\in\mathcal{C}$ as
$n_{c}(k)=\sum_{u\in\mathcal{U}}\mathbb{I}\big(c\in R_{u,k}\big),$ (22)
and the corresponding normalized exposure distribution
$P_{\mathrm{exp},k}(c)=\frac{n_{c}(k)}{\sum_{c^{\prime}\in\mathcal{C}}n_{c^{\prime}}(k)}.$ (23)
The Global Exposure Entropy at cutoff $k$ is then
$\mathrm{Entropy@}k=H_{k}:=-\sum_{c\in\mathcal{C}}P_{\mathrm{exp},k}(c)\,\log_{2}P_{\mathrm{exp},k}(c).$ (24)
Higher values of $H_{k}$ indicate that exposure is distributed more equitably across clusters, while lower values indicate concentration on a small subset of clusters (mode collapse). We use presence in the top-$k$ list (unweighted) to measure exposure concentration; position-weighted variants yield similar trends.
E.2 Yelp Dataset
On the Yelp dataset, we evaluate SPECTRA in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach. In this regime, Logit Probing becomes prohibitively expensive due to the large output space. To address this, we adopt Hierarchical Probing (Algorithm˜3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”). For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”. In inference, SPECTRA first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category. While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
Hierarchical Probing Mapping. We ask Gemini-2.5 Pro and GPT-5 to produce this mapping with the prompt:
You are an expert in ontology construction, taxonomy design, and user preference modeling, with extensive experience in building interpretable, semantically meaningful cluster hierarchies for preference learning. I will provide you a list of raw user preference clusters, each with an associated weight or frequency representing its importance or user interaction volume (e.g., number of views). Bicycles: number of reviews xxx, average ratings xxx;
Massage: number of reviews xxx, average ratings xxx;
Your task is to: 1. Group these raw clusters into high-level categories (Level-1) that are semantically meaningful and interpretable. 2. Assign each raw cluster to one and only one Level-1 category, forming a two-level hierarchy. 3. Return the result as a JSON object, where the keys are Level-1 category names and the values are lists of Level-2 clusters (i.e., raw cluster names). You should aim for: 1. 20 to 26 Level-1 categories total. 2. Coherent semantics within each group. 3. Human-readable, concise names for each Level-1 category (e.g., “Food & Drink”, “Fitness”, “Beauty”).
With the mapping, of which we verified that Level-1 categories are well-balanced, with no single group dominating the total frequency or collapsing into overly broad labels:
L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, SPECTRAs, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
E.3 Real-World Short-Video Dataset
On the short-video dataset, we select the play duration as the proxy metric to represent the user interests. Compared to sparse binary signals like likes or saves, play duration offers a more reliable and verifiable measure of user interest. While some approaches construct weighted combinations of multiple engagement signals, we deliberately opt for a single, interpretable metric to avoid introducing volatility and platform-specific heuristics. The ground-truth distribution is defined as the aggregated play duration across future windows (3, 7, or 14 days). Although we cannot release raw correlation statistics due to data sensitivity and privacy constraints, internal audits confirm strong alignment between play duration and user feedback.
Our analysis targets high-activity users ($\geq$21 active days in the past month), as they represent the platform’s core, most engaged cohort. Since we only have $30$-days of user data, we pick high-activity users for their dense acitivities.
Durations are normalized per user to form a probability distribution over clusters, preventing heavy-usage magnitude from confounding preference shares.
Appendix F Additional Experimental Results
F.1 Evaluation of SPECTRA on Movielens
Figure 4: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions). Prediction windows: long-term and short-term. Metrics: NDCG@$1,5,10$ and JS-Divergence. We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (CPT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model. Note that for the Direct Generation baseline, the JS-divergence will not be available. Sample size: 38,434.
Qwen3-8B DeepSeek-distill-Llama-8B
Prediction Method Entropy@1 Entropy@3 Entropy@5 Entropy@10 Entropy@1 Entropy@3 Entropy@5 Entropy@10
Long A: Direct Generation 1.020 2.446 3.558 3.849 0.140 1.629 2.666 3.447
B: Logit-Probing 1.747 3.130 3.565 3.979 2.307 3.054 3.352 3.813
C: Gen. Class. 1.575 2.732 3.103 3.705 0.952 1.926 2.608 3.329
Short A: Direct Generation .810 2.163 3.581 3.767 0.050 1.635 2.713 3.476
B: Logit-Probing 2.093 3.246 3.612 3.964 2.256 2.922 3.252 3.701
C: Gen. Class. 1.605 2.696 3.211 3.681 0.073 1.972 2.350 3.433
Table 6: Fairness performance on Movielens using the CPT+SFT pipeline. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: entropy score; context window: 3. Best results for each setting are in bold.
Inference procedure of Direct Generation. Unlike SPECTRA, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation .813 .573 .644 .609 .440 .540
B: Logit-Probing .781 .776 .823 .320 .734 .786 .840 .373
C: Gen. Class. .795 .725 .736 .547 .562 .606 .632 .462
Qwen3-Reranker-8B .814 .713 .728 .399
Short A: Direct Generation .701 .506 .571 .499 .369 .450
B: Logit-Probing .667 .675 .740 .466 .542 .619 .714 .514
C: Gen. Class. .638 .661 .789 .538 .499 .490 .501 .551
Qwen3-Reranker-8B .650 .566 .652 .525
3 Long A: Direct Generation .820 .573 .644 .601 .428 .533
B: Logit-Probing .815 .815 .858 .309 .758 .812 .812 .370
C: Gen. Class. .830 .760 .758 .547 .637 .624 .624 .462
Qwen3-Reranker-8B .809 .714 .727 .401
Short A: Direct Generation .702 .495 .554 .492 .361 .453
B: Logit-Probing .686 .674 .707 .460 .501 .612 .712 .507
C: Gen. Class. .661 .631 .671 .539 .492 .499 .507 .545
Qwen3-Reranker-8B .624 .563 .645 .525
5 Long A: Direct Generation .840 .538 .629 .604 .428 .542
B: Logit-Probing .830 .815 .825 .315 .763 .817 .867 .367
C: Gen. Class. .847 .777 .764 .547 .423 .528 .551 .463
Qwen3-Reranker-8B .802 .709 .725 .403
Short A: Direct Generation .723 .496 .546 .493 .359 .454
B: Logit-Probing .716 .720 .781 .460 .496 .606 .708 .507
C: Gen. Class. .673 .635 .676 .535 .489 .498 .507 .546
Qwen3-Reranker-8B .606 .552 .641 .525
8 Long A: Direct Generation .848 .530 .613 .605 .425 .546
B: Logit-Probing .831 .822 .863 .316 .774 .823 .872 .367
C: Gen. Class. .846 .760 .762 .542 .362 .518 .542 .465
Qwen3-Reranker-8B .801 .701 .721 .406
Short A: Direct Generation ..711 .499 .529 .492 .359 .456
B: Logit-Probing .710 .719 .780 .462 .490 .609 .711 .507
C: Gen. Class. .657 .632 .673 .530 .486 .497 .509 .547
Qwen3-Reranker-8B .609 .522 .638 .528
Table 7: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B. Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation .996 .847 .779 .979 .857 .803
B: Logit-Probing .993 .951 .754 .993 .966 .919
C: Gen. Class. .992 .890 .736 .948 .821 .760
Short A: Direct Generation .921 .549 .484 .772 .467 .417
B: Logit-Probing .887 .711 .589 .788 .688 .608
C: Gen. Class. .882 .627 .517 .772 .588 .434
3 Long A: Direct Generation .999 .857 .773 .972 .845 .799
B: Logit-Probing .991 .960 .914 .996 .973 .926
C: Gen. Class. .997 .911 .792 .897 .854 .755
Short A: Direct Generation .926 .555 .488 .775 .466 .429
B: Logit-Probing .909 .738 .617 .755 .693 .618
C: Gen. Class. .902 .654 .518 .775 .601 .446
5 Long A: Direct Generation .998 .866 .767 .976 .842 .804
B: Logit-Probing .996 .961 .913 .993 .973 .931
C: Gen. Class. .998 .906 .789 .869 .857 .756
Short A: Direct Generation .930 .559 .488 .768 .463 .429
B: Logit-Probing .926 .749 .622 .746 .687 .620
C: Gen. Class. .912 .662 .525 .762 .597 .443
8 Long A: Direct Generation .999 .869 .776 .975 .837 .807
B: Logit-Probing .996 .963 .918 .998 .977 .934
C: Gen. Class. .999 .907 .794 .858 .861 .757
Short A: Direct Generation .925 .555 .486 .774 .461 .431
B: Logit-Probing .922 .749 .621 .736 .694 .622
C: Gen. Class. .907 .658 .521 .762 .597 .447
Table 8: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation .072 .299 .547 .071 .301 .563
B: Logit-Probing .072 .342 .637 .073 .350 .658
C: Gen. Class. .072 .317 .552 .068 .289 .533
Short A: Direct Generation .124 .349 .603 .097 .285 .519
B: Logit-Probing .120 .461 .743 .106 .446 .772
C: Gen. Class. .118 .402 .654 .097 .367 .538
3 Long A: Direct Generation .073 .304 .544 .070 .297 .562
B: Logit-Probing .072 .345 .651 .073 .354 .666
C: Gen. Class. .072 .325 .559 .064 .302 .531
Short A: Direct Generation .124 .354 .608 .096 .278 .527
B: Logit-Probing .122 .477 .777 .101 .443 .773
C: Gen. Class. .121 .417 .653 .096 .369 .543
5 Long A: Direct Generation .073 .310 .544 .070 .295 .563
B: Logit-Probing .073 .349 .656 .072 .351 .666
C: Gen. Class. .073 .326 .562 .061 .302 .529
Short A: Direct Generation .125 .357 .609 .095 .278 .529
B: Logit-Probing .125 .486 .784 .099 .441 .780
C: Gen. Class. .123 .423 .662 .095 .368 .542
8 Long A: Direct Generation .073 .309 .547 .070 .292 .564
B: Logit-Probing .073 .348 .655 .073 .353 .668
C: Gen. Class. .073 .324 .562 .060 .304 .529
Short A: Direct Generation .125 .356 .607 .096 .277 .532
B: Logit-Probing .126 .488 .785 .099 .445 .781
C: Gen. Class. .123 .422 .659 .095 .370 .547
Table 9: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold.
(a) Group-level preference evolution as a probability 3D map
(b) Group-level preference evolution as a bar chart
Figure 5: Evolution of group-level movie genre preferences as probability distribution across four time periods. Dataset: Movielens, 125 users over 4 periods. Method: SPECTRA Likelihood-based Probing with CPT+SFT on Qwen3-8B. Note: due to limited space, we only show top-10 clusters (genres) here.
F.2 Evaluation of SPECTRA on Yelp
Qwen3-8B DeepSeek-distill-Llama-8B
Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
Logit Probing .980 .879 .915 .925 .856 .880
Generative Classification .107 .118 .163 .001 .062 .119
L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
Logit Probing .513 .562 .666 .261 .268 .363
Generative Classification .466 .349 .454 .684 .553 .574
Table 10: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task. Results are shown for a context window of 8 using the CPT+SFT pipeline. Best NDCG scores for each level are in bold. L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”. Alternative taxonomies can be easily integrated depending on platform objectives. Sample Size: 5000.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation 0.0015 0.1069 0.2430 0.0127 0.3109 0.3425
B: Logit-Probing 0.8604 0.8436 0.8669 0.6145 0.9581 0.8501 0.8777 0.6499
C: Gen. Class. 0.0945 0.1147 0.1587 0.6969 0.0000 0.1021 0.1437 0.7245
Short A: Direct Generation 0.0149 0.1133 0.2686 0.0127 0.2971 0.3333
B: Logit-Probing 0.7430 0.7467 0.7891 0.6768 0.7072 0.7352 0.7650 0.6969
C: Gen. Class. 0.0398 0.1122 0.1549 0.7301 0.0001 0.0928 0.1468 0.7296
3 Long A: Direct Generation 0.0040 0.1164 0.3381 0.0170 0.3142 0.3450
B: Logit-Probing 0.9557 0.8724 0.9016 0.5976 0.9590 0.8683 0.8866 0.6527
C: Gen. Class. 0.1254 0.1230 0.1684 0.6935 0.0000 0.0647 0.1451 0.7179
Short A: Direct Generation 0.0277 0.1403 0.3556 0.0219 0.3034 0.3379
B: Logit-Probing 0.8743 0.7970 0.8548 0.6611 0.3391 0.4507 0.5194 0.7020
C: Gen. Class. 0.0454 0.1120 0.1577 0.7235 0.0028 0.0552 0.1823 0.7334
5 Long A: Direct Generation 0.0115 0.1279 0.3522 0.0176 0.3136 0.3481
B: Logit-Probing 0.9629 0.8713 0.9075 0.5848 0.7814 0.8057 0.8350 0.6579
C: Gen. Class. 0.1204 0.1243 0.1720 0.6839 0.0004 0.0567 0.1554 0.7143
Short A: Direct Generation 0.1101 0.1472 0.3557 0.0203 0.3001 0.3394
B: Logit-Probing 0.8823 0.7923 0.8585 0.6511 0.6402 0.7237 0.7643 0.6944
C: Gen. Class. 0.0411 0.1137 0.1556 0.7194 0.0905 0.0802 0.1107 0.7317
8 Long A: Direct Generation 0.0081 0.1390 0.3297 0.0156 0.3084 0.3429
B: Logit-Probing 0.9795 0.8787 0.9152 0.5734 0.9246 0.8563 0.8796 0.6546
C: Gen. Class. 0.1070 0.1183 0.1628 0.7012 0.0013 0.0615 0.1187 0.7168
Short A: Direct Generation 0.0739 0.1347 0.3152 0.0176 0.2948 0.3367
B: Logit-Probing 0.8593 0.7874 0.8599 0.6470 0.6809 0.7250 0.7637 0.6984
C: Gen. Class. 0.0402 0.1227 0.1526 0.7341 0.1095 0.0863 0.1292 0.7328
Table 11: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation 0.0054 0.4100 0.2583 0.1423 0.2503 0.2667
B: Logit-Probing 0.9538 0.6749 0.4897 0.9892 0.5025 0.3898
C: Gen. Class. 0.6152 0.3920 0.3456 0.0018 0.3508 0.2656
Short A: Direct Generation 0.0408 0.2252 0.1707 0.0444 0.2130 0.1885
B: Logit-Probing 0.8613 0.3718 0.2754 0.7905 0.3910 0.2628
C: Gen. Class. 0.1453 0.2006 0.1947 0.0006 0.1976 0.1681
3 Long A: Direct Generation 0.0186 0.3880 0.2929 0.1531 0.2516 0.2672
B: Logit-Probing 0.9886 0.6379 0.4957 0.9946 0.5557 0.4017
C: Gen. Class. 0.7713 0.3863 0.3482 0.0006 0.2281 0.2606
Short A: Direct Generation 0.0828 0.2101 0.2060 0.0576 0.2172 0.1902
B: Logit-Probing 0.9526 0.3905 0.3186 0.4106 0.2403 0.2018
C: Gen. Class. 0.1477 0.1963 0.1983 0144 0.1005 0.1634
5 Long A: Direct Generation 0.0690 0.3539 0.2918 0.1609 0.2549 0.2881
B: Logit-Probing 0.9898 0.6293 0.5030 0.9040 0.5651 0.4384
C: Gen. Class. 0.7863 0.4104 0.3592 0.0018 0.1878 0.1997
Short A: Direct Generation 0.3854 0.1894 0.2015 0.0540 0.2172 0.1975
B: Logit-Probing 0.9616 0.3864 0.3268 0.7419 0.4715 0.3124
C: Gen. Class. 0.1339 0.2110 0.1962 0.3025 0.1028 0.0850
8 Long A: Direct Generation 0.0755 0.3220 0.3002 0.1519 0.2533 0.2757
B: Logit-Probing 0.9970 0.6293 0.5070 0.9676 0.5442 0.4164
C: Gen. Class. 0.6783 0.3903 0.3338 0.0078 0.2077 0.2478
Short A: Direct Generation 0.2636 0.1887 0.2048 0.0462 0.2108 0.1954
B: Logit-Probing 0.9652 0.3844 0.3413 0.8025 0.4415 0.2920
C: Gen. Class. 0.1429 0.2394 0.1838 0.3938 0.1122 0.1185
Table 12: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation 0.0007 0.2977 0.3807 0.0173 0.1994 0.4070
B: Logit-Probing 0.1614 0.5224 0.7448 0.1675 0.4038 0.6063
C: Gen. Class. 0.0881 0.2823 0.5051 0.0002 0.2653 0.3910
Short A: Direct Generation 0.0089 0.2678 0.4435 0.0099 0.3070 0.4955
B: Logit-Probing 0.2604 0.5066 0.7148 0.2319 0.5153 0.6654
C: Gen. Class. 0.0304 0.2305 0.4533 0.0002 0.2313 0.3975
3 Long A: Direct Generation 0.0025 0.2850 0.4488 0.0190 0.2032 0.4138
B: Logit-Probing 0.1699 0.4997 0.7660 0.1706 0.4500 0.6296
C: Gen. Class. 0.1149 0.2773 0.5120 0.0001 0.1672 0.3889
Short A: Direct Generation 0.0180 0.2496 0.5500 0.0117 0.3146 0.5038
B: Logit-Probing 0.2899 0.5252 0.8171 0.1288 0.3325 0.5308
C: Gen. Class. 0.0304 0.2214 0.4585 0.0030 0.1171 0.4046
5 Long A: Direct Generation 0.0096 0.2624 0.4490 0.0196 0.2012 0.4378
B: Logit-Probing 0.1666 0.4903 0.7741 0.1532 0.4516 0.6763
C: Gen. Class. 0.1181 0.2950 0.5330 0.0002 0.1340 0.2853
Short A: Direct Generation 0.0926 0.2283 0.5344 0.0113 0.3070 0.5138
B: Logit-Probing 0.2859 0.5151 0.8345 0.2182 0.6206 0.7940
C: Gen. Class. 0.0272 0.2392 0.4568 0.0700 0.1174 0.1942
8 Long A: Direct Generation 0.0105 0.2411 0.4689 0.0183 0.2012 0.4208
B: Logit-Probing 0.1729 0.4966 0.7887 0.1645 0.4377 0.6477
C: Gen. Class. 0.1016 0.2811 0.4921 0.0013 0.1484 0.3565
Short A: Direct Generation 0.0605 0.2223 0.5327 0.0097 0.3088 0.5182
B: Logit-Probing 0.2870 0.5137 0.8638 0.2392 0.5892 0.7530
C: Gen. Class. 0.0301 0.2784 0.4256 0.0908 0.1279 0.2726
Table 13: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold. Level: L1